### v1

# Accelerating Containerized Service Delivery at the Network Edge

Abstract
Efficient container image distribution is crucial for enabling machine learning inference at the network edge, where resource limitations and dynamic network conditions create significant challenges. In this paper, we present PeerSync, a decentralized P2P-based system designed to optimize image distribution in edge environments. PeerSync employs a popularity- and network-aware download engine that dynamically adapts to content popularity and real-time network conditions using a sliding window mechanism. PeerSync further integrates automated tracker election for rapid peer discovery and dynamic cache management for efficient storage utilization. We implement PeerSync with 8000+ lines of Rust code and test its performance extensively on both physical edge devices and Docker-based emulations. Experimental results show that PeerSync delivers a remarkable speed increase of 2.72$\times$, 1.79$\times$, and 1.28$\times$ compared to the Baseline, Dragonfly, and Kraken, respectively, while significantly reducing peak cross-network traffic by 90.72% under congested and varying network conditions.
Index Terms:
edge computing, container image distribution, P2P architecture, model inference.
I Introduction
Model inference at the network edge is becoming increasingly popular in applications such as edge-based video analytics, real-time inference in smart cities, and IoT-based predictive maintenance. To support these scenarios, containerization technologies, such as Docker, are widely used to enable the rapid deployment and management of workloads at scale [38, 44]. Container images, representing multi-layered filesystems, are not merely software artifacts but structured data that must be efficiently stored, distributed, and retrieved across diverse environments, ranging from cloud platforms to the network edge [13, 16, 34, 47, 2, 22].
Deploying a containerized application involves retrieving the necessary image layers from a container registry (e.g., Docker Hub), unpacking them, mounting them using a layered filesystem, and initiating the container entrypoint. These images, structured as a stack of layers representing filesystem segments, are stored and managed in centralized container registries. While these registries efficiently facilitate container image distribution in cloud environments, they encounter unique challenges in edge computing scenarios.
  • High latency. Distributed edge devices face substantial delays when fetching image layers over wide-area networks, leading to degraded responsiveness and increased cold-startup times for services [16, 8, 11].
  • Inefficiency under high concurrency. The non-parallel architecture of container registry uplinks often causes severe bottlenecks when handling concurrent image-pulling requests from multiple edge devices [11]. The growing size of AI models exacerbates this issue.
To address these limitations, peer-to-peer (P2P) architectures have been explored to accelerate container image distribution while alleviating the load on centralized registries [16, 44, 48, 19, 22, 26, 49, 23, 27, 2, 43]. P2P systems inherently offer distributed replication, parallel data transfers, and resilience against single points of failure. Although P2P-based solutions have shown promise in cloud environments, their effectiveness in edge settings is constrained by the unique characteristics of edge computing:
  • Limited network bandwidth. Unlike high-capacity data center networks, edge environments often operate with significantly constrained bandwidth, hindering efficient data exchange [13].
  • Low storage scalability. Portable edge devices, which typically rely on constrained storage media like SD cards or embedded flash memory, often have limited storage. Unlike cloud data centers equipped with technologies like Storage Area Networks [41], their scalability is inherently restricted. Over time, the accumulation of large AI-related images and other data, such as raw sensor outputs, can quickly deplete available storage space.
  • Weak stability under dynamic topology. The dynamic nature of edge networks introduces frequent connectivity changes, requiring robust mechanisms to adaptively manage edge devices and maintain consistency without relying on centralized components such as static trackers [20]. [^1] A tracker is used to identify which peers possess or request a given object in a P2P network. See [20] for more details.
While P2P-based solutions such as EdgePier [8], Kraken [43], and Starlight [11] address some of these issues and achieve great performance in most cases, they are still based on the traditional P2P architecture and consequently fall short in effectively balancing network efficiency, storage optimization, and adaptability to edge-specific conditions. More specifically, they lack fine-grained control mechanisms and rely on static trackers, making them unsuitable for fluctuating network conditions. Starlight [11] minimizes container startup latency by redesigning deployment protocols and storage formats but requires intrusive modifications to cloud registries, limiting its applicability. Moreover, existing solutions often neglect factors such as dynamic network quality, content popularity, and effective storage utilization, leaving room for improvement. For instance, our experiments (detailed in Sec. II) demonstrate that maintaining a single copy of any image layer within a local area network (LAN) can significantly enhance image-fetching speed without overloading local storage.
TABLE I: A comparison between our work and state-of-the-art implementations. ‘F.D.’ stands for ‘Fully Decentralized’, while ‘S.D.’ stands for ‘Semi Decentralized’.
Reference PeerSync Kraken [43] Dragonfly [2] Starlight [11] EdgePier [8] FID [23]
Architecture P2P (F.D.) P2P (S.D.) P2P (S.D.) C/S P2P (F.D.) P2P (S.D.)
P2P flavor Impl. from scratch Impl. from scratch Impl. from scratch N/A IPFS BitTorrent
Scenario Edge Cloud Cloud WAN Edge General
Tracker placement dynamic static static N/A static static
Open-sourced
To overcome these challenges, we propose PeerSync, a non-intrusive, P2P-based system tailored to the unique demands of edge computing. PeerSync constructs a fully decentralized P2P network spanning multiple edge devices across different LANs. It features a P2P image download engine that optimizes distribution by periodically calculating the required content pieces for each missing layer of a requested image, leveraging a scoring system that evaluates both content popularity and the quality of the network connection between peers and the requesting edge device. Unlike traditional P2P architectures that rely on manually configured trackers, PeerSync autonomously elects trackers based on real-time metrics, including bandwidth availability and resource utilization, ensuring resilience and performance. PeerSync also employs a selective deletion mechanism to optimize storage utilization, maintaining only essential image layers within each LAN. We compare PeerSync with state-of-the-art solutions in Table I. In summary, our main contributions are as follows:
  1. 1.
    We accelerate edge container image distribution with a tailored P2P architecture, featuring a high-performance download engine, fault-tolerant tracker with election ability for faster peer discovery, and collaborative dynamic space reclamation for efficient storage management.
  2. 2.
    We dynamically assign scores to peers based on runtime metrics like content popularity and network conditions using a sliding window mechanism, significantly enhancing the flexibility of image distribution in dynamic edge networks.
  3. 3.
    We implement PeerSync with 8000+ lines of Rust code, resulting in a statically linked binary of just 8.8 MB, making it easily deployable on edge devices.
  4. 4.
    We conducted extensive experiments to verify PeerSync’s performance under different network conditions. On average, PeerSync achieves 2.72x faster distribution than the Baseline, 1.79x faster than Dragonfly, and 1.28x faster than Kraken. Additionally, PeerSync reduces peak cross-network traffic by 90.72% under congested and variable network conditions.
II Motivation
In this section, we demonstrate the limitations of existing approaches using Kraken [43] and Starlight [11], motivating the need for a more efficient solution tailored to edge computing environments.
II-A Traditional P2P Approaches
P2P architectures, such as Napster [10], BitTorrent [1], and IPFS [9], provide decentralized mechanisms where nodes act as both consumers and providers. While these systems excel in distributed replication and data sharing, they suffer from a critical drawback: the lack of bandwidth awareness, which results in suboptimal use of network resources, particularly in edge environments.
Figure 1: An edge computing environment with 2 LANs and the corresponding observed results when using BitTorrent for image downloading.
To demonstrate this limitation, we conducted an experiment using Kraken [43]. As depicted on the left side of Fig. 1, our testbed consisted of two LANs connected via a 100 Mbps link. Each LAN included three Raspberry Pi (RPi) 4 Model B devices connected to a 1 Gbps switch. In LAN 1, two hosts seeded a large container image, while requests for the image originated from LAN 2. The experiment was repeated with 1, 2, and 3 local seeding peers in LAN 2. The results, shown in the right side of Fig. 1, highlight a significant inefficiency: despite having local peers available in LAN 2, approximately 10% of the data blocks were still fetched from remote peers in LAN 1, consuming over 95% of the transit network bandwidth. This behavior stems from BitTorrent’s peer selection strategy, which prioritizes global distribution over localized resource optimization. Consequently, the uplink bandwidth became saturated, negating the benefits of local peers and limiting the speed advantage of P2P systems in edge environments.
Tracker-based peer discovery is also vulnerable in dynamic edge environments given its centralized architecture. When a tracker-carrying node fails, it leads to fragmented networks, increased latency, and the requirement of manual intervention [12, 32, 25].
II-B Image Structure-based Approaches
Another class of approaches, exemplified by Starlight [11], attempts to optimize container startup time by analyzing the structures of the requested images. Starlight removes unnecessary components and decouples the download and execution processes to minimize startup latency. While effective for some use cases, these optimizations rely on runtime trace collection, which assumes that future executions will follow similar patterns and the two phases can be effectively parallelized. This creates a critical drawback when dealing with diverse and unpredictable workloads. Moreover, Starlight retains a reliance on centralized registries, limiting its effectiveness in dynamic, resource-constrained edge environments.
Figure 2: Timeline diagram of a large model inference task. Top: Lazy Loading (e.g., Starlight [11]): (1) Download components (PyTorch, cuDNN library, etc.), (2) Download the model, (3) Load the ML framework, (4) Perform inference. Bottom: P2P (e.g., PeerSync): (1) Download the full image, (2) Load the ML framework, (3) Perform inference.
As shown in Fig. 2, Starlight’s lazy loading strategy prioritizes downloading smaller components (e.g., libraries) before larger files like machine learning models. However, this approach provides only minimal speedup for inference tasks, as the critical delay stems from downloading large model files before computation can begin. For example, uncompressed CUDA-accelerated versions of Meta’s Llama 3.1 [15] (e.g., ai/meta-llama:3.1-8B-Instruct-cuda-12.6) require approximately 21.4 GiB of disk space. This image is dominated by two major components: models (15 GiB or 70.10%, consisting of only four files using Safetensors [3]) and frameworks (6.2 GiB or 28.97%, comprising PyTorch [35] and CUDA libraries). Trace-based methods like Starlight struggle with such storage footprints, as the bulk of the data resides in a few large files. In contrast, P2P-based approaches, such as the one proposed in this work, optimize how data is transmitted rather than focusing on when or what to download. By leveraging multiple peers, P2P systems can accelerate the distribution of large data blocks, making them more effective for workloads dominated by large model components.
II-C Insights
Both Kraken and Starlight illustrate shortcomings in existing approaches. Kraken, while leveraging P2P architectures, suffers from bandwidth inefficiencies and excessive reliance on centralized trackers. Starlight, despite its innovative lazy-loading approach, struggles with scalability across diverse model inference workloads and fails to address large model file distribution effectively. These limitations highlight the need for a decentralized, bandwidth-aware solution that maximizes local resource utilization, minimizes bandwidth waste, and adapts dynamically to the constraints of edge environments.
III System Design
III-A Overview
Fig. 3 illustrates the architecture of PeerSync, comprising four main components: Container API Interface (Sec. III-B), Download Engine (Sec. III-C), Embedded Tracker (Sec. III-D), and Cache Manager (Sec. III-E). Container API Interface provides a seamless and non-intrusive replacement for conventional Docker registries by implementing a suite of APIs. Download Engine coordinates the image retrieval process, combining traditional HTTP-based methods with innovative P2P-based approaches. Embedded Tracker maintains an autonomous and self-healing tracking system to ensure swarm integrity, eliminating reliance on centralized components. Cache Manager optimizes storage by managing cached images dynamically, prioritizing frequently used images and removing unused ones with minimal performance impact.
Figure 3: PeerSync architecture. The yellow boxes are PeerSync components.
Fig. 3 also depicts the lifecycle of an image within PeerSync, where red lines represent the retrieval process, and blue lines represent the removal process.
[…]
III-B Container API Interface
Container API Interface leverages the standardized APIs established by the Open Container Initiative (OCI). [^2] https://opencontainers.org/ Specifically, we adapt a subset of the Docker registry APIs to enable comprehensive container image retrieval and management while preserving simplicity for the user. This integration allows for seamless image handling, effectively masking the inherent complexity of the process. A critical aspect of this implementation is the management of image pulls by the Docker engine, which typically involves multiple HTTP requests: one to retrieve the image manifest and several more for each missing layer. The manifest, typically a JSON structure, contains essential details about the requested image, while layer requests return the actual data in the form of tarballs.
Our approach strategically optimizes these two API endpoints:
  • Manifest. Manifests are cached in an in-memory storage structure, which is continuously updated to reflect any upstream changes from Docker Hub.
  • Layer. This endpoint is a streamlined encapsulation that delegates the complexities of the multi-strategy image downloading process to Downloading Engine.
The API encapsulation layer within Docker Engine provides a simplified interface. Upon initialization, PeerSync queries the Docker Engine’s version via a UNIX domain socket and subsequently supplies runtime data to other system components. Essentially, Container API Interface acts as a bridge between Docker Engine and PeerSync’s internal mechanisms. Through this efficient encapsulation, PeerSync maintains precise control and monitoring of the Docker Engine, enabling a focus on optimizing internal operations.
III-C Downloading Engine
Downloading Engine serves as a core component in PeerSync, managing DHT network connections and facilitating popularity- and network-aware P2P downloading upon requests.
III-C1 Request Dispatcher
Request Dispatcher interfaces directly with Container API Interface (Sec. III-B). Upon receiving a download request, it evaluates the optimal strategy for image retrieval, deciding between downloads from a center registry and leveraging the P2P network. The process begins by assessing the size of the requested layer, using information stored in the manifest cache. Next, the Request Dispatcher determines whether the content is available within the nearby P2P network, taking into account time constraints during the search process.
P2P Downloading or Docker Hub? The efficiency of locating peers in the Distributed Hash Table (DHT) network is critical, as this process, which involves multiple network hops, can be time-consuming. Poor network conditions between peers, such as limited bandwidth or packet loss, can further extend the duration of peer discovery. In our context, such delays are unacceptable, as they negatively impact user experience by causing prolonged waiting times. To address this, we implement a time limit for peer aggregation. If image availability is confirmed within this timeframe, it indicates that image holders are present in nearby networks, making P2P downloading the preferred method. However, if the time limit expires without success, it suggests the absence of local image holders or their significant distance from the requesting node, prompting the system to switch to direct downloading from Docker Hub.
TABLE II: Size distribution of popular image layers. Each percentage represents the proportion of layers that are smaller than the specified threshold.
Thres. (B) % $<$ Thres. Thres. (B) % $<$ Thres.
128 1.64 1 Ki 29.27
8 Ki 41.45 512 Ki 47.78
4 Mi 57.38 32 Mi 76.81
256 Mi 97.19 605.73 Mi 100
Partial P2P Downloading. We analyzed the layer sizes of the top 100 popular images from Docker Hub as of July 2024, and as shown in Table II, the majority of these layers are relatively small, with a median size of just 1.03 MiB. Based on this observation, Request Dispatcher employs a partial P2P-based downloading strategy that balances efficiency and resource utilization by avoiding the lengthy processes typically associated with full DHT and tracker-based peer discovery [42]. Comprehensive discovery processes can take disproportionately longer than downloading directly from the central registry. However, relying exclusively on registry downloads for these smaller layers could lead to registry overload. To address this, our partial P2P method prioritizes multicast-based local peer discovery, bypassing the more time-consuming DHT and tracker-based methods. This adjustment significantly reduces both data transfer time and volume. If the requested layer is not found on the local network, Request Dispatcher defaults to direct downloads from Docker Hub. Any subsequent requests for the same layer within the local network are then fulfilled internally, ensuring rapid access while minimizing network congestion on the center registry. This approach effectively optimizes the trade-off between fast access and reduced registry load.
III-C2 Popularity- and Network-Aware Downloading
In PeerSync’s P2P downloading process, content is segmented into blocks, a standard practice that enables simultaneous downloading of different blocks from multiple peers, thereby maximizing bandwidth utilization [46]. When Request Dispatcher opts to download a particular image layer, it provides P2P Downloader with the layer details and relevant peer information, initiating the download process. Fig. 4 illustrates the P2P Downloader’s workflow, which consists of five key stages:
  1. 1.
    The process begins with the selection of a batch of blocks for the current download cycle.
  2. 2.
    Potential peers are sampled based on their dynamically updated scores, maintained by Peer Aggregator. This scoring system assesses peers’ suitability for the download task by considering both content popularity and network quality.
  3. 3.
    The peer with the highest score for each block is then selected to fulfill the download request, ensuring the optimal peer is chosen for each block.
  4. 4.
    A download request is sent to the selected peer.
  5. 5.
    Upon receiving a block, its integrity is verified. If the block fails validation, it is returned to the pending queue for re-download; otherwise, it is stored in the cache and delivered to the client.
Figure 4: Workflow of P2P downloading in PeerSync.
The cycle repeats with a new batch of blocks once the current one is successfully downloaded and validated.
TABLE III: Download times for an 8194.5 MiB image in a 10 Gbps LAN environment with different block sizes.
Block Size (MiB) #Blocks Download Time (s)
256 33 58
128 65 59
32 257 64
16 513 63
8 1025 87
Determining the Block Size. The methodology for determining block size and selecting peers for download requests is crucial to optimizing performance. The block size must be carefully balanced: blocks that are too large reduce parallelism by decreasing the number of blocks available for simultaneous download, while blocks that are too small increase protocol overhead and computational load due to a larger Merkle tree and more hash operations. Our experimental analysis, shown in Table III, used an 8194.5 MiB image in a 10 Gbps LAN environment. The results indicated that smaller block sizes significantly prolonged the download time. Based on these findings, we empirically determined the block size using the following guidelines, where $L_{b}$ denotes the block size and $L_{i}$ denotes the image size:
$L_{b}=\left\{\begin{array}[]{ll}L_{i}/256&\text{if }L_{i}\geq 1024\text{ MiB},\\ L_{i}/64&\text{if 256 MiB }\leq L_{i}<1024\text{ MiB},\\ L_{i}/16&\text{if 16 MiB}\leq L_{i}<256\text{ MiB},\\ L_{i}&\text{otherwise.}\\ \end{array}\right.$ (1)
For smaller images, which can be quickly downloaded from a single source, we consolidate the image into a single block to minimize communication overhead while still applying peer scoring algorithms.
Peer Selection Mechanism. Upon determining the optimal block size, PeerSync proceeds to the critical phase of peer selection. This phase significantly impacts system efficiency, influencing both the download time and the load on network links, particularly in congested conditions. PeerSync’s selection algorithm evaluates peers based on three criteria: network position, network speed, and content popularity. Each peer is assigned a score within the range of [0,100][0,100][ 0 , 100 ], reflecting its suitability for selection, with higher scores correlating to a greater probability of being chosen for content retrieval. In the following, we refer to the device initiating the content request as the client.
Network Position and Speed. The first criterion in PeerSync’s peer selection is network position—whether a peer resides in the same LAN as the client. Local peers are prioritized, receiving the maximum score of 100. For peers outside the LAN, PeerSync employs a speed-based scoring algorithm, where higher network speeds result in higher scores. This strategy favors local peers when available, and otherwise selects the fastest peers. Let $s_{p}^{t}$ denote the speed from peer $p$ to the client at time step $t$, and $\bar{s}^{t}$ the average speed from all peers to the client at time step $t$. Additionally, PeerSync maintains a sliding window $\mathcal{W}_{p}$, of size $L$, for recording historical transmission speeds from peer $p$ to the client. Similarly, $\mathcal{W}$ records the global average speeds. The speed $s_{p}^{t}$ is updated as:
$s_{p}^{t}=\frac{\sum_{s_{p}^{t^{\prime}}\in\mathcal{W}_{p}}s_{p}^{t^{\prime}}\cdot e^{L-t^{\prime}}}{\sum_{s_{p}^{t^{\prime}}\in\mathcal{W}_{p}}e^{L-t^{\prime}}},$ (2)
and the global average speed $\bar{s}^{t}$ is calculated as:
$\bar{s}^{t}=\frac{\sum_{\bar{s}^{t^{\prime}}\in\mathcal{W}}\bar{s}^{t^{\prime}}\cdot e^{L-t^{\prime}}}{\sum_{\bar{s}^{t^{\prime}}\in\mathcal{W}}e^{L-t^{\prime}}}.$ (3)
The network score $\text{net}_{p}^{t}$ of peer $p$ is then obtained as:
$\text{net}_{p}^{t}=s_{p}^{t}-\bar{s}^{t}.$ (4)
After re-scaling, $\text{net}_{p}^{t}$ is mapped to a value within [0,100][0,100][ 0 , 100 ].
Content Popularity. The popularity of an image layer affects its retrieval efficiency. Clients should avoid selecting peers storing rare images to optimize bandwidth usage. Let $\mathcal{P}^{t}$ denote the set of peers known to the client at time step $t$, and $\mathcal{I}_{p}^{t}$ the set of images possessed by peer $p$. We compute the popularity $\rho_{l}$ of layer $l$ as:
$\rho_{l}=\frac{\sum_{p\in\mathcal{P}^{t}}\sum_{i\in\mathcal{I}_{p}^{t}}\epsilon_{l}^{i}}{\sum_{p\in\mathcal{P}^{t}}\sum_{i\in\mathcal{I}_{p}^{t}}1},$ (5)
where $\epsilon_{l}^{i}=0$ if layer $l$ exists in image $i$, and 1 otherwise. The popularity score of peer $p$ is then given by:
$\text{pop}_{p}^{t}=100\times\left(1-\frac{\sum_{i\in\mathcal{I}_{p}^{t}}\sum_{l\in\mathcal{L}_{i}}e^{-\lambda\cdot\rho_{l}}}{\sum_{i\in\mathcal{I}_{p}^{t}}\sum_{l\in\mathcal{L}_{i}}1}\right).$ (6)
Custom Scoring and Softmax Distribution. Recognizing different user needs, PeerSync allows users to define custom scoring functions. The custom score of peer $p$ is denoted by $\text{cst}_{p}^{t}$. The final score (we also name it utility) of peer $p$, denoted by $U(p;t)$, is a weighted sum of network speed, content popularity, and the custom score:
$U(p;t)=\alpha\cdot\text{net}_{p}^{t}+\beta\cdot\text{pop}_{p}^{t}+\gamma\cdot\text{cst}_{p}^{t},$ (7)
where $\alpha$, $\beta$, and $\gamma$ are weights. PeerSync uses a softmax function over the scores to compute the probability distribution for peer selection:
$\Pr\{p;t\}=\frac{\exp(U(p;t))}{\sum_{p^{\prime}\in\mathcal{P}}\exp(U(p^{\prime};t))}.$ (8)
This ensures peers with higher scores are more likely to be selected.
Theoretical Analysis. Let $p_{t}$ be the peer selected by PeerSync and $p_{t}^{*}$ the peer that maximizes utility for a given client at time $t$. The instantaneous regret for selecting peer $p_{t}$ is defined as:
$R_{t}=U(p_{t}^{*};t)-U(p_{t};t).$ (9)
Theorem 1.
PeerSync’s peer selection mechanism achieves sublinear regret in terms of the cumulative difference between the optimal utility and the utility obtained by PeerSync. Specifically, over $T$ rounds of peer selection for different clients and content requests, the cumulative regret is bounded by $O(\sqrt{T})$, i.e.,
$R(T)=\sum_{t=1}^{T}(U(p_{t}^{*};t)-U(p_{t};t))\leq O(\sqrt{T}).$ (10)
The proof is omitted.
Proof.
The regret bound follows from the softmax-based selection strategy employed by PeerSync, where the probability of selecting a peer is a function of its utility score. The softmax function, with a decaying temperature parameter $\tau_{t}$, ensures exploration early on and focuses on exploitation as $\tau_{t}\to 0$, converging to the optimal peer.
The expected instantaneous regret $R_{t}$ is bounded by:
$\mathbb{E}[R_{t}]\leq\frac{C}{\tau_{t}},$ (11)
where $C$ is a constant representing the maximum possible utility difference between peers. Using $\tau_{t}=\frac{\tau_{0}}{\sqrt{t}}$, the cumulative regret over $T$ rounds is bounded by
$R(T)\leq\sum_{t=1}^{T}\frac{C}{\sqrt{t}}=O(\sqrt{T}),$ (12)
which means that PeerSync’s peer selection mechanism has a sublinear regret. ∎
III-D Embedded Autonomous Tracker
Trackers serve as centralized services crucial for facilitating peer discovery within P2P networks [21]. Without a functional tracker, peer discovery times can increase dramatically due to the multi-hop iteration of DHT, leading to inefficiencies. However, the centralized architecture of trackers poses potential points of failure, such as hardware malfunctions, software crashes, network outages, or inadequate maintenance. Additionally, managing a separate tracker infrastructure introduces administrative overhead.
To address these challenges, PeerSync incorporates an embedded autonomous tracker system that dynamically initiates a leader election process when no active trackers are available within the swarm. This system is triggered when a node detects the absence of live trackers or fails to connect to an existing tracker. The autonomous tracker supports the coexistence of multiple trackers but activates its election mechanism when all trackers are unavailable, particularly in local swarm regions. The node that identifies this situation initiates the election process using the FloodMax algorithm. [^3] FloodMax is a well-known leader election protocol in distributed systems. See [28] for more details. This algorithm enables each node to broadcast its stability metrics, such as uptime, to neighboring nodes. To mitigate the high communication overhead inherent in the FloodMax algorithm for large-scale networks, PeerSync implements path-pruning techniques [33]. These techniques reduce unnecessary message traffic, especially across network boundaries, achieving near-linear message complexity. This optimization minimizes resource consumption while ensuring the timely election of a new tracker.
Integrated directly into PeerSync’s codebase, Embedded Autonomous Tracker eliminates the need for manual configuration, significantly reducing the risk of single points of failure. By enhancing resilience and adaptability, this feature ensures uninterrupted operations in dynamic and distributed environments.
III-E Cache Manager
Efficient storage management is essential for maintaining the stability of low-end edge devices, which often operate with constrained resources [29]. Over time, PeerSync accumulates residual images in its cache, consuming valuable disk space, even when the associated applications are no longer active. Furthermore, upper-level applications require spare storage for data-heavy tasks, such as storing raw sensor metrics and datasets. Edge devices, typically reliant on flash-based storage without advanced virtualization technologies like Ceph [45], face additional limitations. Consequently, PeerSync must maintain its storage footprint as small as possible while preserving performance.
PeerSync addresses these challenges with a dynamic cache cleaner that leverages usage history to manage storage. This component monitors image usage through the Docker Engine, gathering statistics to determine which images to remove when storage reaches critical levels (e.g., less than 10% free space or a user-defined threshold). Image removal is prioritized based on three factors: last access time, image size, and content popularity (we consider both local and global popularity).
PeerSync’s storage management algorithm extends the classic Least Recently Used (LRU) cache replacement strategy by incorporating an additional metric: cache miss cost. Unlike traditional LRU, which assumes uniform memory access cost across an address space [39, 40], PeerSync accounts for the non-trivial time differences in retrieving an image:
  • Images present within the same LAN incur minimal retrieval cost, making them candidates for removal if redundancy exists.
  • Images that are the only copy within a LAN but have multiple copies elsewhere are prioritized for removal based on the number of external replicas.
  • Sole known copies of images are given the lowest removal priority, ensuring critical data remains accessible unless absolutely necessary.
This popularity-aware algorithm enables PeerSync to balance storage efficiency and performance, optimizing both retrieval times and disk utilization.
III-F Implementation Structure
PeerSync is designed with a message-passing architecture. Each component operates as an asynchronous task within the host system, with well-defined inter-component interfaces enabling communication and interaction. This approach provides both reliability and extensibility. In terms of reliability, isolated components ensure that failures do not affect the entire system. Any failed component can be quickly restarted to maintain continuous service. Extensibility is achieved through modular design, enabling easy updates to incorporate new features, adhere to evolving specifications, and support additional image formats.
PeerSync supports multiple peer discovery methods that function independently. This ensures Downloading Engine can immediately initiate fetching when an image is available from the P2P swarm, similar to BitTorrent. If P2P fetching is not feasible, PeerSync seamlessly falls back to direct HTTP downloading, ensuring uninterrupted operation.
TABLE IV: The container images chosen for evaluation.
Name Tag Service Compressed Size Description
redhat/granite-3-1b-a400m-instruct latest NLP 1.47 GB 1B finetuned IBM Granite 3.0 [18]
ai/meta-llama 3.1-8B-Instruct NLP 14.91 GB Llama 3.1 model by Meta [15]
cvisionai/segment-anything latest Vision 5.2 GB Segment Anything model by Meta [24]
langchain/langchain latest NLP 437.57 MB LLM application adaption framework[5]
pytorch/pytorch 2.5.1-cuda12.4-cudnn9-runtime General 3.11 GB Deep learning framework[35]
tensorflow/tensorflow nightly-gpu General 3.61 GB Deep learning framework[7]
IV Evaluation
We conduct extensive experimental evaluations in two environments: a Docker Compose-based emulation on a high-performance x86_64 host (Intel Xeon Silver 4214) and physical edge devices (Raspberry Pi 4 Model B). The emulation enables the simulation of large-scale environments, while the physical setup validates the practical feasibility of PeerSync in real-world scenarios. We select a range of popular AL/ML and commonly used container images of varying sizes, from small base images to large language model (LLM) containers, as listed in Table IV. These images represent diverse use cases, allowing us to evaluate system performance under varying workloads and conditions.
We evaluated PeerSync against two popular container image distribution systems: Dragonfly [2] and Kraken [43], as well as the conventional HTTP-based pull method, referred to as the Baseline. Dragonfly, hosted by the Cloud Native Computing Foundation (CNCF) as an Incubating Level Project, [^4] https://d7y.io/ leverages peer-to-peer (P2P) technology to accelerate file distribution and container image delivery. It is designed to enhance efficiency in distributed environments by reducing reliance on centralized registries. Kraken, an open-source project developed by Uber and deployed in production since early 2018, [^5] https://github.com/uber/kraken also employs P2P technology for Docker image management, replication, and distribution. Kraken is particularly tailored for hybrid cloud environments, offering decentralized capabilities to optimize large-scale image deployments. Although Starlight [11] also focuses on improving container image delivery, it relies on fundamentally different mechanisms, such as runtime trace collection and dependence on centralized registries. These differences make direct comparisons with PeerSync less relevant (see Sec. II). Consequently, we did not include Starlight in our comparisons.
Figure 5: Average AI/ML image distribution time under various network conditions (log scale). The $x$-axis is the parameter $A$. The missing points mean that the download operation could not be completed during the limited time frame.
IV-A Docker Compose-based Emulation
IV-A1 Experimental Setup
We create 10 Linux bridge networks ($n$}, where $n=1$ to 101010) to emulate a distributed environment with interconnected routers. Each network contained a router for inter-network communication and seven worker nodes requesting container images. The routers were configured with bandwidth limits (50 Mbps to 1 Gbps), latency, and packet loss using tc(8) [6], simulating real-world ISP and transit networks. Intra-network bandwidth was set to 1 Gbps with zero packet loss.
To model real-world conditions, we vary model inference request frequencies based on the following equation:
$t_{i}\sim\text{Poisson}^{-1}\left(\text{random}\left(0.001,A\cdot e^{\frac{B}{\text{s}_{i}}}\right)\right),$
where $\text{s}_{i}$ is the image size, and $A$ and $B$ are tunable parameters. Higher values of $A$ and $B$ increased the frequency of image-pulling requests, simulating high-demand scenarios. Background traffic generated by iPerf [4] created additional congestion, reflecting the shared link usage typical in real-world applications. All centralized components, including registries, Dragonfly databases, and Kraken trackers, were placed in net_worker1 to simulate centralized cloud environments.
IV-A2 Image Distribution Time
TABLE V: Average Speedup under high request frequency compared to the Baseline (smaller is better).
Net. Profile Baseline Dragonfly[2] Kraken[43] PeerSync
Stable 100% 75.01% 80.13% 77.85%
Congested 100% 64.27% 61.89% 45.89%
Varying 100% 65.63% 46.96% 36.71%
Fig. 5 illustrates the average container image distribution times under different network conditions, providing critical insights into the importance of local data prioritization. Since the Baseline method shows significant performance degradation as request frequency increases, a logarithmic scale (base 222) is used on the $y$-axis for clarity.
Stable network conditions (see top of Fig. 5). In environments with sufficient and stable network resources, all P2P-based methods, including PeerSync, exhibit similar performance trends as request frequency increases (see Table V). While distribution latency increases slightly with higher request loads, P2P methods consistently outperform the Baseline HTTP approach. This aligns with the known benefits of P2P models, such as efficient data sharing and reduced reliance on centralized registries, which have been extensively validated in data center contexts. The trend is particularly pronounced for large LLM images: the average distribution time roughly quadruples as request frequency increases, highlighting the critical role of P2P technologies in facilitating the deployment of large modern AI/ML models with billions of parameters.
Congested network conditions (see lower-left of Fig. 5). In environments with bandwidth limitations, packet loss, and high latency, the Baseline method suffers severe performance degradation due to the long-tail effects of centralized pulling. Remote data transfers are particularly affected, often exceeding the default time limit of 1200 seconds, as indicated by the missing points in the green Baseline lines. In contrast, PeerSync and other P2P-based methods maintain better performance by distributing the load across multiple peers, effectively mitigating bandwidth constraints. Notably, Dragonfly shows a sharp performance drop for larger images due to its reliance on centralized components such as schedulers and managers. As communication between edge nodes and these components becomes unreliable, Dragonfly’s ability to coordinate cloud-edge interactions weakens, resulting in bottlenecks during image serving. PeerSync, by contrast, leverages its decentralized architecture to achieve consistent performance, even under congestion.
Variable network conditions (see lower-right of Fig. 5). This scenario simulates real-world edge environments where latency, packet loss, and bandwidth fluctuate significantly, and participating nodes frequently join and leave the network. Such instability disrupts inter-peer connections and hinders image fetching, even for smaller images like langchain/langchain. Compared to other P2P-based solutions, PeerSync demonstrates superior resilience under these conditions. By prioritizing local peer synchronization and minimizing dependence on centralized components, PeerSync reduces reliance on specific, potentially unstable links and dynamically explores all available links to maximize performance. This strategy not only improves distribution efficiency but also enhances fault tolerance, which is critical in environments with unpredictable network behavior.
PeerSync’s ability to optimize local storage and peer communication makes it particularly robust in dynamic edge environments. As shown in Table V, PeerSync outperforms the Baseline, Dragonfly, and Kraken under unstable, high-frequency scenarios by average factors of 2.72, 1.79, and 1.28, respectively. [^6] In our experiments, the time limit for each download request was set to 1200 seconds. In some instances, the Baseline was unable to finish within this time frame, resulting in missing data points when calculating the average speedup. We assigned a value of 1200 seconds to the missing data points. Consequently, the actual speedup relative to the Baseline is higher than the calculated value. The speedup is even more pronounced for larger container images, reinforcing PeerSync’s effectiveness in addressing the challenges of large-scale image distribution.
IV-A3 Uplink Occupancy
Tables VI, VII, and VIII illustrate the cross-network traffic for each image distribution method across three distinct network profiles: stable, congested, and unstable conditions. Cross-network traffic represents data that traverses routers and encounters bandwidth limits, latency, and packet loss, effectively simulating the challenges of data traversing upstream ISP transit networks and the Internet. Unlike local network bandwidth, which is often abundant, cross-network bandwidth is typically constrained by external factors such as physical link limitations and tariff plans. Thus, minimizing cross-network traffic is critical, particularly in edge computing, where the goal is to prioritize intra-LAN resources as much as possible.
TABLE VI: Sum of cross-network traffic under sufficient and stable network conditions.
Solution Maximum (Gbps) Average (Gbps)
PeerSync 5.07 0.82
Kraken [43] 6.61 0.93
Dragonfly [2] 8.80 1.37
Baseline 11.91 6.18
Sufficient and stable network conditions. As shown in Table VI, PeerSync consistently minimizes cross-network bandwidth usage under stable conditions. By prioritizing local peer data, PeerSync achieves an average bandwidth consumption of only 0.82 Gbps, significantly lower than Kraken (0.93 Gbps), Dragonfly (1.37 Gbps), and the Baseline (6.18 Gbps). While Dragonfly and Kraken leverage P2P techniques, their reliance on centralized components for coordination and scheduling results in higher cross-network traffic compared to PeerSync. The Baseline method incurs the highest bandwidth usage due to its centralized pulling mechanism, which does not take advantage of local data sharing.
TABLE VII: Sum of cross-network traffic under congested but stable network conditions.
Solution Maximum (Gbps) Average (Gbps)
PeerSync 11.50 0.91
Kraken [43] 13.76 1.18
Dragonfly [2] 11.94 4.69
Baseline 10.36 9.81
Congested but stable network conditions. In congested but stable network conditions (Table VII), PeerSync continues to outperform other methods by maintaining an average cross-network traffic of just 0.91 Gbps. The Baseline approach shows the lowest maximum bandwidth (10.36 Gbps), which is a result of its slower and less efficient data retrieval. However, it saturates network capacity with an average bandwidth usage of 9.81 Gbps, underscoring its inefficiency in managing congested networks. Kraken and Dragonfly perform better than the Baseline but still exhibit performance degradation due to their centralized dependencies. Dragonfly, in particular, suffers from its reliance on schedulers and managers, resulting in an average traffic of 4.69 Gbps, over five times higher than PeerSync.
TABLE VIII: Sum of cross-network traffic under congested and unstable (i.e. changing packet loss, latency, and bandwidth) network conditions.
Solution Maximum (Gbps) Average (Gbps)
PeerSync 9.15 0.76
Kraken [43] 10.99 0.89
Dragonfly [2] 10.13 4.01
Baseline 8.87 6.34
Congested and unstable network conditions. Table VIII highlights the performance of each method under congested and unstable network conditions, characterized by fluctuating packet loss, latency, and bandwidth. PeerSync demonstrates remarkable robustness, maintaining the lowest average cross-network traffic at 0.76 Gbps. Kraken performs slightly worse, with an average traffic of 0.89 Gbps, but still manages to leverage its decentralized architecture to reduce reliance on external networks. Dragonfly, on the other hand, suffers significant performance degradation due to its dependence on centralized coordination, resulting in an average traffic of 4.01 Gbps. The Baseline method experiences the lowest maximum bandwidth usage (8.87 Gbps) but continues to exhibit high average traffic (6.34 Gbps), reflecting its inability to adapt to varying network conditions. This highlights the inefficiency of centralized pulling in handling unstable and dynamic environments.
The results across all network profiles clearly demonstrate PeerSync’s superior ability to minimize cross-network bandwidth consumption. This efficiency is achieved by leveraging local resources and reducing reliance on external network traffic, which is particularly critical in edge environments where bandwidth is constrained and often unstable.
TABLE IX: Number of edge devices in a LAN and the average image distribution time (over 100 image retrieval requests) provided by LAN cache.
Number of edge devices within a LAN 1 2 3 4 5 6 7 8 9 10
Average image distribution time (s) 9.11 8.33 9.69 11.34 8.87 8.06 5.08 2.86 2.89 2.19
TABLE X: Number of nodes (1 $\sim$ 5) and comparison of the sum of cache space occupied by Cache Cleaner and the LRU policy.
Number of edge devices within a LAN 1 2 3 4 5 6 7 8 9 10
Space occupied by Cache Cleaner (MiB) 196 388 475 589 593 668 905 1259 1199 1420
Space occupied by LRU (MiB) 110 206 411 701 837 948 973 1164 1056 1701
Figure 6: Average image distribution time under congested and unstable (i.e. changing packet loss, latency, and bandwidth) network conditions (log scale). The $x$-axis is the parameter $A$. (Smaller images)
IV-A4 Cache Strategy
We evaluated the Cache Cleaner algorithm by analyzing its effectiveness in managing requests and accelerating image retrieval within a LAN. Table IX shows the relationship between the number of edge devices in a LAN and the average time to download images. Initially, as the number of nodes increases, image distribution times rise slightly due to simultaneous requests to the upstream registry. However, as the LAN becomes more populated, the local cache network is increasingly leveraged, resulting in significantly reduced average retrieval times.
As nodes in the LAN form a collaborative cache, the Cache Cleaner algorithm effectively retains images that are likely to benefit the entire network. This approach optimizes local data sharing and minimizes redundant requests to external networks. The results demonstrate that as the cache network scales, the average download time decreases dramatically, underscoring the efficiency of PeerSync’s collaborative caching mechanism.
Comparison with the LRU policy. To further validate the Cache Cleaner’s performance, we compared its total cache usage against the Least Recently Used (LRU) policy in Table X. Cache Cleaner consistently achieves better space utilization by coordinating cache management across neighboring peers. This approach avoids redundant caching of identical content across multiple nodes, enabling more efficient storage utilization within the LAN. In contrast, LRU operates independently on each node and lacks collaboration, leading to higher total cache consumption. Cache Cleaner’s ability to prioritize fast intra-LAN retrieval not only optimizes storage but also improves overall system efficiency.
Distribution of smaller images. Although our primary focus is on large AI/ML images, smaller, non-AI base images are also widely used in practical setups. To emulate the operation of Docker Hub, which serves popular images in large volumes daily, we selected the 10 most downloaded images from Docker Hub and increased the request frequency by adjusting the parameter $A$. Fig. 6 presents the results, demonstrating PeerSync’s ability to maintain its performance advantage over other P2P-based solutions, even under high-frequency request conditions for smaller images. These results highlight PeerSync’s capability to handle the significant volumes typical of Docker Hub operations, proving its versatility across diverse workloads.
The experiment results validate the efficiency of PeerSync’s Cache Cleaner algorithm in optimizing local caching, reducing redundant external requests, and improving retrieval times within a LAN. Its collaborative caching strategy outperforms traditional policies like LRU, offering better space utilization and accelerated data distribution. Additionally, PeerSync demonstrates its ability to efficiently handle high-frequency requests for smaller images, further reinforcing its robustness and scalability across varied workloads and network conditions. This makes PeerSync an efficient solution for both edge and cloud-based environments.
IV-B Real-World Experiments
We deployed a physical testbed using Raspberry Pi devices to compare PeerSync’s performance against the Baseline, Dragonfly, and Kraken. As shown in Fig. 7, the setup comprises six Raspberry Pi (RPi) 4 Model B devices connected via two layer 2 switches, each with a 1 Gbps link speed. Each LAN contains three RPis, with the two switches connected to a common router. The router is configured to forward packets between the networks while limiting the inter-LAN bandwidth to 100 Mbps. To simplify deployment, static routes are configured on the router. This setup is similar to the Docker Compose-based emulation described earlier, with two key differences: a) Physical hardware and real-world networking scenarios are used; b) The aarch64 version of PeerSync is deployed instead of the amd64 version to accommodate the ARM-based Raspberry Pi architecture.
Figure 7: The testbed for real-world experiments.
Benchmark requests were made to agents running on each RPi to measure image distribution times. Table XI presents the 90th and 100th percentiles of distribution times achieved by each system. The results demonstrate that PeerSync significantly outperforms the other systems in real-world conditions.
  • At the 90th percentile (P90), PeerSync achieves an image distribution time of 190.79 seconds, representing a 39.7% improvement over Kraken, a 60.9% improvement over Dragonfly, and a 75.9% improvement over the Baseline.
  • At the 99th percentile (P99), PeerSync maintains its advantage, with a distribution time of 243.89 seconds, which is 27.8% faster than Kraken, 53.9% faster than Dragonfly, and 69.9% faster than the Baseline.
TABLE XI: The 90th and 99th percentile of image distribution times.
Solution Baseline Dragonfly Kraken PeerSync
P90 (s) 790.13 487.50 316.42 190.79
P99 (s) 811.23 528.49 337.75 243.89
The results show that PeerSync performs better than the other methods significantly in our real-world experiments.
V Related Works
Numerous solutions have been explored to enhance distribution speeds in container image delivery. In this section, we briefly review key developments in this area, covering both container-specific solutions and methodologies applied to other resource types.
Before the era of edge computing, several pioneering works focused on fast resource delivery in data center networks. VDN [36] leverages hierarchical network topology and shared chunks between virtual machine images to achieve a 30–80$\times$ speedup for large virtual machine images under heavy traffic in data center networks. Similarly, VMTorrent [37] improves P2P methods used in live streaming, employing block prioritization, profile-based execution prefetch, on-demand fetch, and decoupling of virtual machine image presentation from the underlying data stream. Experiments show VMTorrent can achieve a 30x speedup over traditional network storage.
Since the rise of Docker, containers have become central to deploying workloads in data centers, and distributing container images has emerged as a key issue. While similar to virtual machine images, containers have unique properties, such as a layered structure and registry-based hosting. DevOps practices demand more automated and integrated workload deployment solutions. FID [22], an early P2P-based approach for fast container image distribution, adapts BitTorrent to provide an integrated distribution system. In industry, Dragonfly [2] and Kraken [43] are prominent P2P container image distribution systems. These methods outperform traditional approaches by utilizing spare bandwidth between clients.
However, these solutions do not explicitly address edge computing environments. While Dragonfly [2] and Kraken [43] are P2P-based, they rely on multiple centralized components, which can degrade service quality in the event of component failure. Edge environments are less stable than traditional cloud facilities, and deploying these centralized components in the cloud requires manual configuration across different edge locations. Additionally, edge devices typically have weaker network capabilities, so unrestricted use of P2P protocols can lead to uplink congestion. EdgePier [8], based on IPFS [9], provides a fully decentralized container image distribution system tailored to edge computing. Similarly, Gazzetti et al. [17] proposed a decentralized solution with managers that compute optimal network topologies.
Beyond P2P, another trend involves distributed storage to form a shared storage layer among participating nodes. CoMICon [31], Cider [14], and Wharf [49] use distributed filesystems to accelerate container provisioning. However, these methods are primarily focused on data centers, and implementing distributed storage in edge environments remains challenging [30].
More recently, Starlight [11] has introduced a novel approach by redesigning the container image architecture. It transmits only the necessary components, significantly reducing transfer times. However, Starlight’s reliance on specific container engines limits its applicability, and it continues to use a traditional client-server registry architecture, which restricts throughput under high bandwidth loads. Future work could integrate Starlight with PeerSync to achieve (a) distributed transmission of large data chunks and (b) selective transmission of necessary data, further improving the containerization ecosystem.
VI Concluding Remarks
In this paper, we introduced PeerSync, an innovative fully decentralized image distribution system tailored for containerized model inference workloads at the network edge. PeerSync leverages advanced P2P downloading techniques, outperforming traditional methods by dynamically adapting to fluctuating network conditions and content popularity. Its embedded autonomous tracker enhances resilience by eliminating single points of failure, making it particularly effective in edge environments characterized by unstable networks and resource-constrained devices. Additionally, the cache cleaner efficiently optimizes storage usage while maintaining performance, addressing a key limitation of edge devices.
Future work offers several opportunities to extend PeerSync’s capabilities. a) Dynamic block sizing: While the current static block size approach performs well, incorporating adaptive block sizes with multiple Merkle tree versions could further enhance efficiency with minimal overhead. b) Support for heterogeneous LANs: Adapting to diverse network profiles, such as WLANs, using a combination of user-defined tags and auto-detection techniques could improve performance in non-uniform environments. c) Plugin-based algorithm design: Providing users with greater flexibility through plugin-based customization of algorithms could further expand PeerSync’s adaptability.
Integrating PeerSync with Starlight [11] represents another exciting avenue for exploration. The combination of PeerSync’s decentralized, adaptive framework and Starlight’s innovative protocols, filesystem, and image storage formats could pave the way for next-generation containerization ecosystems, optimized for the evolving model inference demands at the network edge.

### v2

@@
-Accelerating Containerized Service Delivery at the Network Edge
+PeerSync: Accelerating Containerized Model Inference at the Network Edge
 Abstract
-Efficient container image distribution is crucial for enabling machine learning inference at the network edge, where resource limitations and dynamic network conditions create significant challenges.
-In this paper, we present PeerSync, a decentralized P2P-based system designed to optimize image distribution in edge environments.
-PeerSync employs a popularity- and network-aware download engine that dynamically adapts to content popularity and real-time network conditions using a sliding window mechanism.
-PeerSync further integrates automated tracker election for rapid peer discovery and dynamic cache management for efficient storage utilization.
-We implement PeerSync with 8000+ lines of Rust code and test its performance extensively on both physical edge devices and Docker-based emulations.
-Experimental results show that PeerSync delivers a remarkable speed increase of 2.72$\times$, 1.79$\times$, and 1.28$\times$ compared to the Baseline, Dragonfly, and Kraken, respectively, while significantly reducing peak cross-network traffic by 90.72% under congested and varying network conditions.
+[…]
 Index Terms:
-edge computing, container image distribution, P2P architecture, model inference.
+Edge computing, container image distribution, P2P architecture, local area network, model inference.
 I Introduction
@@
 These images, structured as a stack of layers representing filesystem segments, are stored and managed in centralized container registries.
-While these registries efficiently facilitate container image distribution in cloud environments, they encounter unique challenges in edge computing scenarios.
-  • High latency.
-Distributed edge devices face substantial delays when fetching image layers over wide-area networks, leading to degraded responsiveness and increased cold-startup times for services [16, 8, 11].
-  • Inefficiency under high concurrency.
-The non-parallel architecture of container registry uplinks often causes severe bottlenecks when handling concurrent image-pulling requests from multiple edge devices [11].
+While these registries efficiently facilitate container image distribution in cloud environments, they encounter unique challenges in edge computing scenarios: (i) High latency: Distributed edge devices face substantial delays when fetching image layers over wide-area networks, leading to degraded responsiveness and increased cold-startup times for services [4, 9, 10]; (ii) Inefficiency under high concurrency: The non-parallel architecture of container registry uplinks often causes severe bottlenecks when handling concurrent image-pulling requests from multiple edge devices [10].
 The growing size of AI models exacerbates this issue.
-To address these limitations, peer-to-peer (P2P) architectures have been explored to accelerate container image distribution while alleviating the load on centralized registries [16, 44, 48, 19, 22, 26, 49, 23, 27, 2, 43].
+To address these limitations, peer-to-peer (P2P) architectures have been explored to accelerate container image distribution while alleviating the load on centralized registries [4, 2, 11, 12, 13, 14, 8, 15, 7, 16].
 P2P systems inherently offer distributed replication, parallel data transfers, and resilience against single points of failure.
 Although P2P-based solutions have shown promise in cloud environments, their effectiveness in edge settings is constrained by the unique characteristics of edge computing:
-  • Limited network bandwidth. Unlike high-capacity data center networks, edge environments often operate with significantly constrained bandwidth, hindering efficient data exchange [13].
+  • Limited network bandwidth. Unlike high-capacity data center networks, edge environments often has significantly constrained bandwidth, hindering fast data exchange [3].
+  • Dynamic network topology.
+The dynamic nature of edge networks introduces frequent connectivity changes, requiring robust mechanisms to adaptively manage edge devices and maintain consistency without relying on centralized components such as static trackers [17].
   • Low storage scalability.
@@
 Over time, the accumulation of large AI-related images and other data, such as raw sensor outputs, can quickly deplete available storage space.
-  • Weak stability under dynamic topology.
-The dynamic nature of edge networks introduces frequent connectivity changes, requiring robust mechanisms to adaptively manage edge devices and maintain consistency without relying on centralized components such as static trackers [20].
-[^1] A tracker is used to identify which peers possess or request a given object in a P2P network.
-See [20] for more details.
+Table I: Comparison between PeerSync and state-of-the-art works. ‘F.D.’: ‘Fully Decentralized’, ‘S.D.’: ‘Semi Decentralized’.
+Work PeerSync Kraken [16] Dragonfly [7] Starlight [10] EdgePier [9] FID [8]
+Architecture P2P (F.D.) P2P (S.D.) P2P (S.D.) C/S P2P (F.D.) P2P (S.D.)
+P2P flavor Impl. from scratch Impl. from scratch Impl. from scratch - IPFS BitTorrent
+Scenario General Cloud Cloud WAN Edge General
+Tracker placement dynamic static static - static static
 While P2P-based solutions such as EdgePier [9], Kraken [16], and Starlight [10] address some of these issues and achieve great performance in most cases, they are still based on the traditional P2P architecture and consequently fall short in effectively balancing network efficiency, storage optimization, and adaptability to edge-specific conditions.
@@
 Moreover, existing solutions often neglect factors such as dynamic network quality, content popularity, and effective storage utilization, leaving room for improvement.
-For instance, our experiments (detailed in Sec.
+For instance, our preliminary experiments (detailed in Sec.
 II) demonstrate that maintaining a single copy of any image layer within a local area network (LAN) can significantly enhance image-fetching speed without overloading local storage.
-TABLE I: A comparison between our work and state-of-the-art implementations. ‘F.D.’ stands for ‘Fully Decentralized’, while ‘S.D.’ stands for ‘Semi Decentralized’.
-Reference PeerSync Kraken [43] Dragonfly [2] Starlight [11] EdgePier [8] FID [23]
-Architecture P2P (F.D.) P2P (S.D.) P2P (S.D.) C/S P2P (F.D.) P2P (S.D.)
-P2P flavor Impl. from scratch Impl. from scratch Impl. from scratch N/A IPFS BitTorrent
-Scenario Edge Cloud Cloud WAN Edge General
-Tracker placement dynamic static static N/A static static
-Open-sourced
-To overcome these challenges, we propose PeerSync, a non-intrusive, P2P-based system tailored to the unique demands of edge computing.
-PeerSync constructs a fully decentralized P2P network spanning multiple edge devices across different LANs.
+Crucially, the edge is not a minimized data center.
+Systems like Kraken [16] and Dragonfly [7], though highly optimized for homogeneous, stable cloud networks, assume persistent centralized trackers or super-nodes that become single points of failure under edge churn.
+EdgePier [9] improves decentralization but still presumes stable anchor nodes, which rarely exist in mobile or intermittently connected edge deployments.
+Meanwhile, proposals leveraging IPFS or BitTorrent [8] inherit protocol overheads ill-suited for low-resource devices.
+Even more recent learning-based approaches [19, 20, 21, 22], while theoretically appealing, require GPU/NPU resources, extensive training data, and stable feedback loops, rendering them impractical on typical edge hardware such as Raspberry Pi.
+To overcome these challenges, we propose PeerSync, a non-intrusive, P2P-based system tailored to the unique demands of the edge.
+PeerSync constructs a fully decentralized P2P network across different LANs.
 It features a P2P image download engine that optimizes distribution by periodically calculating the required content pieces for each missing layer of a requested image, leveraging a scoring system that evaluates both content popularity and the quality of the network connection between peers and the requesting edge device.
@@
   1. 1.
-    We accelerate edge container image distribution with a tailored P2P architecture, featuring a high-performance download engine, fault-tolerant tracker with election ability for faster peer discovery, and collaborative dynamic space reclamation for efficient storage management.
+    We design an edge container image distribution system PeerSync, featuring a high-performance downloading engine, fault-tolerant tracker with self-healing ability, and collaborative dynamic space reclamation for efficient caching.
   2. 2.
-    We dynamically assign scores to peers based on runtime metrics like content popularity and network conditions using a sliding window mechanism, significantly enhancing the flexibility of image distribution in dynamic edge networks.
+    PeerSync dynamically assigns scores to peers.
+Its scoring function jointly optimizes for (i) local network proximity (e.g., same subnet), (ii) real-time bandwidth stability, and (iii) layer popularity.
+This ensures that the majority of traffic remains within the LAN, directly addressing the high-latency and limited-bandwidth constraints of WAN links.
   3. 3.
     We implement PeerSync with 8000+ lines of Rust code, resulting in a statically linked binary of just 8.8 MB, making it easily deployable on edge devices.
+PeerSync is compatible with the Open Container Initiative (OCI) standard and integrates transparently with existing container runtimes.
   4. 4.
-    We conducted extensive experiments to verify PeerSync’s performance under different network conditions.
-On average, PeerSync achieves 2.72x faster distribution than the Baseline, 1.79x faster than Dragonfly, and 1.28x faster than Kraken.
-Additionally, PeerSync reduces peak cross-network traffic by 90.72% under congested and variable network conditions.
+[…]
 II Motivation
-In this section, we demonstrate the limitations of existing approaches using Kraken [43] and Starlight [11], motivating the need for a more efficient solution tailored to edge computing environments.
+In this section, we empirically and analytically dissect the limitations of two representative approaches: Kraken [16] and Starlight [10], motivating the need for a more efficient solution tailored to edge computing environments.
 II-A Traditional P2P Approaches
-P2P architectures, such as Napster [10], BitTorrent [1], and IPFS [9], provide decentralized mechanisms where nodes act as both consumers and providers.
-While these systems excel in distributed replication and data sharing, they suffer from a critical drawback: the lack of bandwidth awareness, which results in suboptimal use of network resources, particularly in edge environments.
+P2P architectures such as Napster [23], BitTorrent [24], and IPFS [25] enable decentralized content sharing by allowing nodes to act as both consumers and providers.
+While these systems excel in cloud or data center settings with abundant bandwidth and stable connectivity, they suffer from a critical flaw in edge environments: the lack of bandwidth awareness and network topology sensitivity.
+This leads to inefficient utilization of scarce inter-LAN links and undermines the very benefits P2P promises.
 Figure 1: An edge computing environment with 2 LANs and the corresponding observed results when using BitTorrent for image downloading.
-To demonstrate this limitation, we conducted an experiment using Kraken [43].
-As depicted on the left side of Fig.
-1, our testbed consisted of two LANs connected via a 100 Mbps link.
-Each LAN included three Raspberry Pi (RPi) 4 Model B devices connected to a 1 Gbps switch.
-In LAN 1, two hosts seeded a large container image, while requests for the image originated from LAN 2.
-The experiment was repeated with 1, 2, and 3 local seeding peers in LAN 2.
-The results, shown in the right side of Fig.
-1, highlight a significant inefficiency: despite having local peers available in LAN 2, approximately 10% of the data blocks were still fetched from remote peers in LAN 1, consuming over 95% of the transit network bandwidth.
-This behavior stems from BitTorrent’s peer selection strategy, which prioritizes global distribution over localized resource optimization.
-Consequently, the uplink bandwidth became saturated, negating the benefits of local peers and limiting the speed advantage of P2P systems in edge environments.
-Tracker-based peer discovery is also vulnerable in dynamic edge environments given its centralized architecture.
-When a tracker-carrying node fails, it leads to fragmented networks, increased latency, and the requirement of manual intervention [12, 32, 25].
+To illustrate this, we conducted an experiment using Kraken [16], which builds upon BitTorrent’s protocol.
+As shown on the left side of Fig.
+1, our testbed comprised two LANs connected via a 100 Mbps link, representative of typical edge-to-edge or edge-to-cloud uplinks.
+Each LAN included three Raspberry Pi 4 Model B devices (each with 4 GB RAM and 32 GB eMMC storage) connected through a 1 Gbps switch.
+In LAN 1, two hosts served as seeders for a large container image; download requests originated from LAN 2, with 1, 2, and 3 local peers progressively added as potential sources.
+The results (right side of Fig.
+1) reveal a striking inefficiency: even when local peers were available in LAN 2, Kraken still fetched approximately 10% of the image blocks from remote peers in LAN 1.
+Crucially, this small fraction consumed over 95% of the inter-LAN bandwidth.
+This behavior stems from BitTorrent’s peer selection strategy, which prioritizes global peer diversity and piece availability over network locality.
+As a result, the narrow uplink becomes saturated, delaying downloads for all nodes and negating the latency advantage of local caching.
+Beyond bandwidth waste, P2P systems like Kraken impose significant storage pressure on edge devices.
+In a separate experiment, we distributed the top-10 most popular Docker Hub images across Kraken nodes.
+After downloading and decompressing these images, each node consumed 1408.54 MiB of disk space.
+For resource-constrained edge devices (e.g., those using SD cards or embedded flash with limited write endurance), such storage overhead is unsustainable, especially when images accumulate over time alongside sensor data or model checkpoints.
+Furthermore, Kraken’s reliance on a centralized tracker for peer discovery introduces a single point of failure.
+In dynamic edge environments, where devices frequently join, leave, or experience intermittent connectivity, tracker outages lead to network fragmentation, prolonged peer discovery latency, and often require manual intervention to restore service [26, 27, 28].
+This centralized coordination model is fundamentally at odds with the autonomy and resilience required in edge deployments.
 II-B Image Structure-based Approaches
-Another class of approaches, exemplified by Starlight [11], attempts to optimize container startup time by analyzing the structures of the requested images.
-Starlight removes unnecessary components and decouples the download and execution processes to minimize startup latency.
-While effective for some use cases, these optimizations rely on runtime trace collection, which assumes that future executions will follow similar patterns and the two phases can be effectively parallelized.
-This creates a critical drawback when dealing with diverse and unpredictable workloads.
-Moreover, Starlight retains a reliance on centralized registries, limiting its effectiveness in dynamic, resource-constrained edge environments.
+An alternative strategy, exemplified by Starlight [10], seeks to minimize container startup latency by analyzing image structures and runtime behavior.
+Starlight removes unnecessary components (a process known as debloating [29, 30]) and decouples the download and execution phases through lazy loading: it prioritizes “hot” files (e.g., libraries) based on runtime traces collected during a profiling run.
+While effective for certain applications, this approach faces three critical limitations in edge AI scenarios:
+  • Starlight assumes workload homogeneity and predictability.
+Starlight requires users to convert images into a custom format and execute them once to collect file-access traces.
+This profiling step assumes that future executions will follow similar patterns, which breaks down when deploying diverse or evolving AI workloads (e.g., switching from object detection to LLM inference).
+At scale, maintaining accurate traces for hundreds of edge applications becomes impractical.
+  • Starlight offers minimal benefit for large-model inference tasks.
+As shown in Fig.
+2, inference pipelines are typically dominated by two phases: (1) fetching the model and framework, and (2) computation.
+Crucially, the entire model must be available before inference can begin, as models are often stored as a few large files (e.g., in Safetensors [31] or PyTorch format [32]).
+Consider Meta’s Llama 3.1 [33] (ai/meta-llama:3.1-8B-Instruct-cuda-12.6).
+When uncompressed, this image occupies approximately 21.4 GiB of disk space, composed of only two major parts: Model weights: 15 GiB (70.10%), stored in just 4 files using Safetensors [31]; ML framework and dependencies: 6.2 GiB (28.97%), including PyTorch [32] and CUDA libraries.
+In such cases, downloading libraries first provides negligible speedup since the critical path remains the transfer of the 15 GiB model file.
+Trace-based methods like Starlight cannot accelerate this bottleneck because they optimize what and when to download, not how fast the data can be fetched.
 Figure 2: Timeline diagram of a large model inference task.
-Top: Lazy Loading (e.g., Starlight [11]): (1) Download components (PyTorch, cuDNN library, etc.), (2) Download the model, (3) Load the ML framework, (4) Perform inference.
-Bottom: P2P (e.g., PeerSync): (1) Download the full image, (2) Load the ML framework, (3) Perform inference.
-As shown in Fig.
-2, Starlight’s lazy loading strategy prioritizes downloading smaller components (e.g., libraries) before larger files like machine learning models.
-However, this approach provides only minimal speedup for inference tasks, as the critical delay stems from downloading large model files before computation can begin.
-For example, uncompressed CUDA-accelerated versions of Meta’s Llama 3.1 [15] (e.g., ai/meta-llama:3.1-8B-Instruct-cuda-12.6) require approximately 21.4 GiB of disk space.
-This image is dominated by two major components: models (15 GiB or 70.10%, consisting of only four files using Safetensors [3]) and frameworks (6.2 GiB or 28.97%, comprising PyTorch [35] and CUDA libraries).
-Trace-based methods like Starlight struggle with such storage footprints, as the bulk of the data resides in a few large files.
-In contrast, P2P-based approaches, such as the one proposed in this work, optimize how data is transmitted rather than focusing on when or what to download.
-By leveraging multiple peers, P2P systems can accelerate the distribution of large data blocks, making them more effective for workloads dominated by large model components.
+Top: Lazy Loading (e.g., Starlight [10]): Download components (PyTorch, cuDNN library, etc.) $\to$ Download the model $\to$ Load the ML framework $\to$ Perform inference.
+Bottom: P2P (e.g., PeerSync): Download the full image $\to$ Load the ML framework $\to$ Perform inference.
+In addition, Starlight remains dependent on centralized registries.
+Despite its optimizations, it still pulls base images from Docker Hub or private registries, inheriting all the latency, bandwidth, and availability challenges of cloud-centric architectures.
+In disconnected or bandwidth-constrained edge settings, this dependency severely limits deployability.
 II-C Insights
@@
 These limitations highlight the need for a decentralized, bandwidth-aware solution that maximizes local resource utilization, minimizes bandwidth waste, and adapts dynamically to the constraints of edge environments.
-III System Design
-III-A Overview
-Fig.
-3 illustrates the architecture of PeerSync, comprising four main components: Container API Interface (Sec.
-III-B), Download Engine (Sec.
-III-C), Embedded Tracker (Sec.
-III-D), and Cache Manager (Sec.
-III-E).
-Container API Interface provides a seamless and non-intrusive replacement for conventional Docker registries by implementing a suite of APIs.
-Download Engine coordinates the image retrieval process, combining traditional HTTP-based methods with innovative P2P-based approaches.
-Embedded Tracker maintains an autonomous and self-healing tracking system to ensure swarm integrity, eliminating reliance on centralized components.
-Cache Manager optimizes storage by managing cached images dynamically, prioritizing frequently used images and removing unused ones with minimal performance impact.
-Figure 3: PeerSync architecture. The yellow boxes are PeerSync components.
-Fig. 3 also depicts the lifecycle of an image within PeerSync, where red lines represent the retrieval process, and blue lines represent the removal process.
-[…]
-[…]
-III-B Container API Interface
-Container API Interface leverages the standardized APIs established by the Open Container Initiative (OCI).
-[^2] https://opencontainers.org/ Specifically, we adapt a subset of the Docker registry APIs to enable comprehensive container image retrieval and management while preserving simplicity for the user.
-This integration allows for seamless image handling, effectively masking the inherent complexity of the process.
-A critical aspect of this implementation is the management of image pulls by the Docker engine, which typically involves multiple HTTP requests: one to retrieve the image manifest and several more for each missing layer.
-The manifest, typically a JSON structure, contains essential details about the requested image, while layer requests return the actual data in the form of tarballs.
-Our approach strategically optimizes these two API endpoints:
-  • Manifest. Manifests are cached in an in-memory storage structure, which is continuously updated to reflect any upstream changes from Docker Hub.
-  • Layer. This endpoint is a streamlined encapsulation that delegates the complexities of the multi-strategy image downloading process to Downloading Engine.
-The API encapsulation layer within Docker Engine provides a simplified interface.
-Upon initialization, PeerSync queries the Docker Engine’s version via a UNIX domain socket and subsequently supplies runtime data to other system components.
-Essentially, Container API Interface acts as a bridge between Docker Engine and PeerSync’s internal mechanisms.
-Through this efficient encapsulation, PeerSync maintains precise control and monitoring of the Docker Engine, enabling a focus on optimizing internal operations.
-III-C Downloading Engine
-Downloading Engine serves as a core component in PeerSync, managing DHT network connections and facilitating popularity- and network-aware P2P downloading upon requests.
-III-C1 Request Dispatcher
-Request Dispatcher interfaces directly with Container API Interface (Sec.
-III-B).
-Upon receiving a download request, it evaluates the optimal strategy for image retrieval, deciding between downloads from a center registry and leveraging the P2P network.
-The process begins by assessing the size of the requested layer, using information stored in the manifest cache.
-Next, the Request Dispatcher determines whether the content is available within the nearby P2P network, taking into account time constraints during the search process.
-P2P Downloading or Docker Hub?
-The efficiency of locating peers in the Distributed Hash Table (DHT) network is critical, as this process, which involves multiple network hops, can be time-consuming.
-Poor network conditions between peers, such as limited bandwidth or packet loss, can further extend the duration of peer discovery.
-In our context, such delays are unacceptable, as they negatively impact user experience by causing prolonged waiting times.
-To address this, we implement a time limit for peer aggregation.
-If image availability is confirmed within this timeframe, it indicates that image holders are present in nearby networks, making P2P downloading the preferred method.
-However, if the time limit expires without success, it suggests the absence of local image holders or their significant distance from the requesting node, prompting the system to switch to direct downloading from Docker Hub.
-TABLE II: Size distribution of popular image layers. Each percentage represents the proportion of layers that are smaller than the specified threshold.
-Thres. (B) % $<$ Thres. Thres. (B) % $<$ Thres.
+III The PeerSync System
+III-A Problem Formulation
+We consider a distributed environment where nodes (including cloud instances and edge devices) pull container images from a central registry.
+Each image consists of a manifest and multiple immutable layers.
+When many nodes concurrently pull the same image, the registry becomes a bottleneck, leading to high latency and WAN bandwidth consumption.
+Let $\mathcal{P}$ be the set of peers, each holding a subset of layers $\mathcal{L}_{p}\subset\mathcal{L}$, where $p\in\mathcal{P}$.
+Peers are connected via heterogeneous, dynamic network links with unknown bandwidth and availability.
+A requesting peer $p_{\text{req}}$ must fetch all layers of a target image as quickly as possible, without prior knowledge of which peers have which layers or their current upload capacity.
+The problem is to minimize the end-to-end image pull time by opportunistically leveraging P2P transfers among peers, while operating without global coordination, runtime modifications, or assumptions about peer stability or connectivity.
+III-B System Architecture Overview
+PeerSync is designed as a drop-in enhancement for existing container ecosystems.
+As illustrated in Fig.
+3, the system comprises four core components: API Interface, Downloading Engine, Embedded Tracker, and Cache Manager, which are orchestrated along two orthogonal threads: (i) the pull path, which handles on-demand layer retrieval, and (ii) the cache maintenance path, which runs continuously in the background to manage storage and peer metadata.
+The entire system exposes a standard OCI Distribution API, making it fully compatible with Docker, Podman, and other OCI-compliant clients.
+No modifications to the container runtime or user workflow are required; peers simply configure their client to use PeerSync as the registry endpoint.
+Figure 3: The architecture of PeerSync.
+A typical pull request begins when a container runtime issues an HTTP GET for a layer.
+The request is intercepted by the API Interface, which checks local cache availability.
+If the layer is absent, the request is forwarded to the Downloading Engine, where a lightweight Request Dispatcher decides whether to fetch the layer from upstream (via HTTP) or initiate a P2P transfer based on multiple factors.
+When P2P is selected, the engine queries the Embedded Tracker for candidate peers and coordinates block-level downloads.
+Upon successful retrieval, the layer is stored in the local Cache, and the response is returned to the client.
+Concurrently, the Cache Manager monitors global access patterns across layers and performs eviction based on a cost-aware policy that considers layer size, replication count within the LAN, and reuse likelihood.
+The Embedded Tracker, operating independently, maintains an up-to-date view of nearby peers through periodic gossip and self-election, ensuring coordination remains decentralized and resilient.
+PeerSync attempts to join the P2P swarm during application startup.
+During bootstrap, PeerSync reads predefined nodes from the configuration and broadcasts bootstrap messages locally to join the swarm.
+If the bootstrap phase fails, PeerSync runs in a degraded mode as a proxy for the upstream registry and continues to bootstrap until a full P2P connection can be established.
+The predefined node list is dynamically updated with high-uptime peers, enabling PeerSync to withstand environmental changes.
+III-C Component Design
+III-C1 Container API Interface
+Modern container ecosystems achieve remarkable interoperability through adherence to the OCI standards.
+The OCI defines both an Image Specification (governing image format and manifest structure) and a Distribution Specification (defining HTTP-based registry APIs for pulling images).
+To ensure PeerSync operates as a drop-in replacement without requiring any modification to existing runtimes or user workflows, it must fully comply with these standards.
+Crucially, the container runtime initiates image pulls via a well-defined sequence: one request for the image manifest (a JSON document listing layer digests and metadata), followed by one request per missing layer (each returning a gzipped tarball).
+Any deviation from this protocol would break compatibility.
+Thus, the primary motivation is to provide a transparent, standards-compliant facade that hides the internal P2P complexity while preserving the exact API contract expected by the runtime.
+The Container API Interface implements a subset of the OCI Distribution Specification. It serves as the sole external entry point for the container runtime.
+  • Manifest Handling.
+The manifest is typically small (often <5<5 KB) but frequently accessed and subject to upstream updates (e.g., latest tag).
+To ensure low-latency responses and immediate consistency, manifests are stored in an in-memory cache.
+This cache is kept up-to-date by periodically polling the upstream registry or by invalidating entries upon detecting a new pull request for a potentially stale tag.
+  • Layer Handling.
+Layer requests are not served directly.
+Instead, the interface acts as a lightweight dispatcher, forwarding each blob request (identified by its digest) to the Downloading Engine (Sec.
+III-C).
+The engine is responsible for fulfilling the request, whether from local cache, a LAN peer, or the upstream registry, and streaming the data back through this interface.
+Figure 4: The state machine of P2P downloading in PeerSync.
+III-C2 Downloading Engine
+A naive P2P approach, which attempts to find peers for every layer, fails to account for the empirical reality of container image composition.
+As shown in Table II, nearly half of all layers are smaller than 512 KiB, with a median size of just 1.03 MiB.
+For such small payloads, the latency of peer discovery in a Distributed Hash Table (DHT) or even a local multicast can exceed the time required to download the layer directly from a nearby registry.
+Conversely, large layers (e.g., base OS images) can benefit immensely from parallel, multi-source P2P downloads.
+The engine’s core motivation is to make a per-layer, context-aware decision that optimizes for both speed and system stability, dynamically choosing between P2P and direct upstream fetching.
+Table II: Layer size distribution of Docker Hub Top 100 images. Each percentage represents the proportion of layers that are smaller than the specified threshold.
+Thres. (B) % << Thres. Thres. (B) % << Thres.
 128 1.64 1 Ki 29.27
@@
 4 Mi 57.38 32 Mi 76.81
-256 Mi 97.19 605.73 Mi 100
-Partial P2P Downloading.
-We analyzed the layer sizes of the top 100 popular images from Docker Hub as of July 2024, and as shown in Table II, the majority of these layers are relatively small, with a median size of just 1.03 MiB.
-Based on this observation, Request Dispatcher employs a partial P2P-based downloading strategy that balances efficiency and resource utilization by avoiding the lengthy processes typically associated with full DHT and tracker-based peer discovery [42].
-Comprehensive discovery processes can take disproportionately longer than downloading directly from the central registry.
-However, relying exclusively on registry downloads for these smaller layers could lead to registry overload.
-To address this, our partial P2P method prioritizes multicast-based local peer discovery, bypassing the more time-consuming DHT and tracker-based methods.
-This adjustment significantly reduces both data transfer time and volume.
-If the requested layer is not found on the local network, Request Dispatcher defaults to direct downloads from Docker Hub.
-Any subsequent requests for the same layer within the local network are then fulfilled internally, ensuring rapid access while minimizing network congestion on the center registry.
-This approach effectively optimizes the trade-off between fast access and reduced registry load.
-III-C2 Popularity- and Network-Aware Downloading
-In PeerSync’s P2P downloading process, content is segmented into blocks, a standard practice that enables simultaneous downloading of different blocks from multiple peers, thereby maximizing bandwidth utilization [46].
-When Request Dispatcher opts to download a particular image layer, it provides P2P Downloader with the layer details and relevant peer information, initiating the download process.
-Fig.
-4 illustrates the P2P Downloader’s workflow, which consists of five key stages:
-  1. 1.
-    The process begins with the selection of a batch of blocks for the current download cycle.
-  2. 2.
-    Potential peers are sampled based on their dynamically updated scores, maintained by Peer Aggregator.
-This scoring system assesses peers’ suitability for the download task by considering both content popularity and network quality.
-  3. 3.
-    The peer with the highest score for each block is then selected to fulfill the download request, ensuring the optimal peer is chosen for each block.
-  4. 4.
-    A download request is sent to the selected peer.
-  5. 5.
-    Upon receiving a block, its integrity is verified.
-If the block fails validation, it is returned to the pending queue for re-download; otherwise, it is stored in the cache and delivered to the client.
+256 Mi 97.19 605.73 Mi 100.00
+The Downloading Engine is structured around a Request Dispatcher and a P2P Downloader.
+Request Dispatcher
+Upon receiving a layer digest from the API interface, the dispatcher first checks with the local Cache Manager (Sec.
+III-C4) and serves the content if locally present.
+Otherwise, a three-stage decision process is executed to retrieve the content from external sources.
+(i) It first queries the other nodes in the same LAN and fetches from the neighbor nodes.
+(ii) If not local, the dispatcher consults the manifest cache to obtain the layer’s size $L_{i}$ to determine if a P2P process shall be started.
+If $L_{i}$ is below a configurable threshold $\theta$ (e.g., 1 MiB), the request is routed directly to the upstream registry.
+This is justified by the data in Table II, which shows that a policy of direct fetch for small layers can handle a large fraction of requests with minimal latency.
+(iii) For layers larger than $\theta$, the dispatcher initiates a P2P discovery process.
+However, this process is bounded by a dynamic timeout $\tau$, defined as the 95th percentile of recent round-trip times (RTTs) over a 10-second sliding window.
+If no suitable peer is found within $\tau$, the system falls back to the upstream registry to avoid indefinite user wait times.
+This logic can be viewed as a state machine (Fig.
+4), but is now implemented as a streamlined, asynchronous workflow to reduce overhead.
+Popularity- and Network-Aware P2P Downloader
+PeerSync segments each image layer into fixed-size blocks to enable concurrent downloads from multiple peers.
+Upon receiving a download request from the Request Dispatcher, the P2P Downloader initiates a five-stage workflow (Fig.
+5): (i) select a batch of blocks for the current cycle; (ii) sample candidate peers based on dynamically updated scores maintained by the Peer Aggregator; (iii) assign each block to the highest-scoring peer; (iv) issue download requests; and (v) validate received blocks via cryptographic hash.
+During this phase, PeerSync re-queues failed blocks or caching and delivers valid ones.
 Figure 5: Workflow of P2P downloading in PeerSync.
-The cycle repeats with a new batch of blocks once the current one is successfully downloaded and validated.
 Table III: Download times for an 8194.5 MiB image in a 10 Gbps LAN environment with different block sizes.
@@
 Determining the block size.
-The methodology for determining block size and selecting peers for download requests is crucial to optimizing performance.
-The block size must be carefully balanced: blocks that are too large reduce parallelism by decreasing the number of blocks available for simultaneous download, while blocks that are too small increase protocol overhead and computational load due to a larger Merkle tree and more hash operations.
-Our experimental analysis, shown in Table III, used an 8194.5 MiB image in a 10 Gbps LAN environment.
-The results indicated that smaller block sizes significantly prolonged the download time.
-Based on these findings, we empirically determined the block size using the following guidelines, where $L_{b}$ denotes the block size and $L_{i}$ denotes the image size:
-$L_{b}=\left\{\begin{array}[]{ll}L_{i}/256&\text{if }L_{i}\geq 1024\text{ MiB},\\ L_{i}/64&\text{if 256 MiB }\leq L_{i}<1024\text{ MiB},\\ L_{i}/16&\text{if 16 MiB}\leq L_{i}<256\text{ MiB},\\ L_{i}&\text{otherwise.}\\ \end{array}\right.$ (1)
-For smaller images, which can be quickly downloaded from a single source, we consolidate the image into a single block to minimize communication overhead while still applying peer scoring algorithms.
-Peer Selection Mechanism.
-Upon determining the optimal block size, PeerSync proceeds to the critical phase of peer selection.
-This phase significantly impacts system efficiency, influencing both the download time and the load on network links, particularly in congested conditions.
-PeerSync’s selection algorithm evaluates peers based on three criteria: network position, network speed, and content popularity.
-Each peer is assigned a score within the range of [0,100][0,100][ 0 , 100 ], reflecting its suitability for selection, with higher scores correlating to a greater probability of being chosen for content retrieval.
-In the following, we refer to the device initiating the content request as the client.
-Network Position and Speed.
-The first criterion in PeerSync’s peer selection is network position—whether a peer resides in the same LAN as the client.
-Local peers are prioritized, receiving the maximum score of 100.
-For peers outside the LAN, PeerSync employs a speed-based scoring algorithm, where higher network speeds result in higher scores.
-This strategy favors local peers when available, and otherwise selects the fastest peers.
-Let $s_{p}^{t}$ denote the speed from peer $p$ to the client at time step $t$, and $\bar{s}^{t}$ the average speed from all peers to the client at time step $t$.
-Additionally, PeerSync maintains a sliding window $\mathcal{W}_{p}$, of size $L$, for recording historical transmission speeds from peer $p$ to the client.
-Similarly, $\mathcal{W}$ records the global average speeds.
-The speed $s_{p}^{t}$ is updated as:
+Block size critically balances parallelism and overhead.
+Excessively large blocks limit concurrency; overly small ones inflate Merkle tree depth and hash computation costs.
+Empirical evaluation using an 8194.5 MiB image in a 10 Gbps LAN (Table III) shows optimal performance at 16-128 MiB.
+Accordingly, PeerSync sets block size $L_{b}$ adaptively based on image size $L_{i}$:
+[…]
+For tiny images whose size $L_{i}<16$ MiB, a single block is used to minimize coordination overhead.
+Peer selection rules. Peer selection combines three criteria into a unified utility score $U(p;t)\in[0,100]$: network position content popularity, and customized logic.
+  1. 1.
+    Network position.
+Peer selection uses a network-aware score that favors peers with higher effective throughput.
+LAN-local peers receive the maximum score of 100.
+For remote peers, the score is based on observed download speed, not low-level metrics like packet loss, as congestion control is handled by the OS kernel [34, 35, 36, 37]; using throughput avoids interfering with kernel mechanisms and better reflects end-to-end performance.
+Each peer $p$ maintains a sliding window $\mathcal{W}_{p}$ of past speeds $\{s_{p}^{t^{\prime}}\}$.
+The current speed estimate $s_{p}^{t}$ is an exponentially weighted average that prioritizes recent samples:
     $s_{p}^{t}=\frac{\sum_{s_{p}^{t^{\prime}}\in\mathcal{W}_{p}}s_{p}^{t^{\prime}}\cdot e^{L-t^{\prime}}}{\sum_{s_{p}^{t^{\prime}}\in\mathcal{W}_{p}}e^{L-t^{\prime}}},$ (2)
-and the global average speed $\bar{s}^{t}$ is calculated as:
+    where $L$ is the current logical time. A global baseline $\bar{s}^{t}$ is computed similarly over all recent transfers:
     $\bar{s}^{t}=\frac{\sum_{\bar{s}^{t^{\prime}}\in\mathcal{W}}\bar{s}^{t^{\prime}}\cdot e^{L-t^{\prime}}}{\sum_{\bar{s}^{t^{\prime}}\in\mathcal{W}}e^{L-t^{\prime}}}.$ (3)
-The network score $\text{net}_{p}^{t}$ of peer $p$ is then obtained as:
-$\text{net}_{p}^{t}=s_{p}^{t}-\bar{s}^{t}.$ (4)
-After re-scaling, $\text{net}_{p}^{t}$ is mapped to a value within [0,100][0,100][ 0 , 100 ].
-Content Popularity.
-The popularity of an image layer affects its retrieval efficiency.
-Clients should avoid selecting peers storing rare images to optimize bandwidth usage.
-Let $\mathcal{P}^{t}$ denote the set of peers known to the client at time step $t$, and $\mathcal{I}_{p}^{t}$ the set of images possessed by peer $p$.
-We compute the popularity $\rho_{l}$ of layer $l$ as:
-$\rho_{l}=\frac{\sum_{p\in\mathcal{P}^{t}}\sum_{i\in\mathcal{I}_{p}^{t}}\epsilon_{l}^{i}}{\sum_{p\in\mathcal{P}^{t}}\sum_{i\in\mathcal{I}_{p}^{t}}1},$ (5)
-where $\epsilon_{l}^{i}=0$ if layer $l$ exists in image $i$, and 1 otherwise. The popularity score of peer $p$ is then given by:
-$\text{pop}_{p}^{t}=100\times\left(1-\frac{\sum_{i\in\mathcal{I}_{p}^{t}}\sum_{l\in\mathcal{L}_{i}}e^{-\lambda\cdot\rho_{l}}}{\sum_{i\in\mathcal{I}_{p}^{t}}\sum_{l\in\mathcal{L}_{i}}1}\right).$ (6)
-Custom Scoring and Softmax Distribution.
-Recognizing different user needs, PeerSync allows users to define custom scoring functions.
-The custom score of peer $p$ is denoted by $\text{cst}_{p}^{t}$.
-The final score (we also name it utility) of peer $p$, denoted by $U(p;t)$, is a weighted sum of network speed, content popularity, and the custom score:
+    The raw network advantage is $\text{net}_{p}^{t}=s_{p}^{t}-\bar{s}^{t}$, which is then linearly rescaled to $[0,100]$ (with negatives clamped to the minimum known score) to yield the final network score.
+This relative scoring adapts to heterogeneous and dynamic network conditions.
+  2. 2.
+    Content popularity. Content popularity discourages reliance on peers holding rare layers. Let $\rho_{l}\in[0,1]$ denote the fraction of known image instances containing layer $l$:
+    $\rho_{l}=\frac{\sum_{p\in\mathcal{P}^{t}}\sum_{i\in\mathcal{I}_{p}^{t}}\epsilon_{l}^{i}}{\sum_{p\in\mathcal{P}^{t}}\sum_{i\in\mathcal{I}_{p}^{t}}1},\quad\epsilon_{l}^{i}=\begin{cases}1&\text{if }l\in i,\\ 0&\text{otherwise}.\end{cases}$ (4)
+    Peers periodically exchange layer sets via differential updates; unchanged states are acknowledged with sequence numbers to minimize overhead. The popularity score for peer $p$ is then:
+    $\text{pop}_{p}^{t}=100\times\left(1-\frac{\sum_{i\in\mathcal{I}_{p}^{t}}\sum_{l\in\mathcal{L}_{i}}e^{-\lambda\cdot\rho_{l}}}{\sum_{i\in\mathcal{I}_{p}^{t}}\sum_{l\in\mathcal{L}_{i}}1}\right),$ (5)
+    which down-weights peers storing rarer content.
+The rationale is to preserve the bandwidth of these critical peers, ensuring they remain available to serve the rare content that only they possess, thus enhancing the overall availability of all layers in the swarm.
+  3. 3.
+    Extensibility via custom scoring.
+To accommodate the diverse and heterogeneous nature of edge deployments, PeerSync’s scoring function is designed with an extensibility hook for domain-specific policies, represented by a custom score $\text{cst}_{p}^{t}$.
+This architectural choice allows administrators to inject logic tailored to their unique operational constraints without modifying the core system.
+For instance, in a wireless edge environment, an administrator may wish to deprioritize peers with weak signal strength to conserve their battery and bandwidth.
+This can be achieved by defining the custom score $\text{cst}_{p}^{t}$ as a function of the peer’s real-time signal strength, thereby penalizing nodes with poor connectivity in the selection process.
+The total utility is a weighted sum:
 $U(p;t)=\alpha\cdot\text{net}_{p}^{t}+\beta\cdot\text{pop}_{p}^{t}+\gamma\cdot\text{cst}_{p}^{t},$ (6)
-where $\alpha$, $\beta$, and $\gamma$ are weights. PeerSync uses a softmax function over the scores to compute the probability distribution for peer selection:
-$\Pr\{p;t\}=\frac{\exp(U(p;t))}{\sum_{p^{\prime}\in\mathcal{P}}\exp(U(p^{\prime};t))}.$ (8)
-This ensures peers with higher scores are more likely to be selected.
-Theoretical Analysis.
-Let $p_{t}$ be the peer selected by PeerSync and $p_{t}^{*}$ the peer that maximizes utility for a given client at time $t$.
-The instantaneous regret for selecting peer $p_{t}$ is defined as:
-$R_{t}=U(p_{t}^{*};t)-U(p_{t};t).$ (9)
-Theorem 1.
-PeerSync’s peer selection mechanism achieves sublinear regret in terms of the cumulative difference between the optimal utility and the utility obtained by PeerSync.
-Specifically, over $T$ rounds of peer selection for different clients and content requests, the cumulative regret is bounded by $O(\sqrt{T})$, i.e.,
-$R(T)=\sum_{t=1}^{T}(U(p_{t}^{*};t)-U(p_{t};t))\leq O(\sqrt{T}).$ (10)
-The proof is omitted.
-Proof.
-The regret bound follows from the softmax-based selection strategy employed by PeerSync, where the probability of selecting a peer is a function of its utility score.
-The softmax function, with a decaying temperature parameter $\tau_{t}$, ensures exploration early on and focuses on exploitation as $\tau_{t}\to 0$, converging to the optimal peer.
-The expected instantaneous regret $R_{t}$ is bounded by:
-$\mathbb{E}[R_{t}]\leq\frac{C}{\tau_{t}},$ (11)
-where $C$ is a constant representing the maximum possible utility difference between peers. Using $\tau_{t}=\frac{\tau_{0}}{\sqrt{t}}$, the cumulative regret over $T$ rounds is bounded by
-$R(T)\leq\sum_{t=1}^{T}\frac{C}{\sqrt{t}}=O(\sqrt{T}),$ (12)
-which means that PeerSync’s peer selection mechanism has a sublinear regret. ∎
-III-D Embedded Autonomous Tracker
-Trackers serve as centralized services crucial for facilitating peer discovery within P2P networks [21].
-Without a functional tracker, peer discovery times can increase dramatically due to the multi-hop iteration of DHT, leading to inefficiencies.
-However, the centralized architecture of trackers poses potential points of failure, such as hardware malfunctions, software crashes, network outages, or inadequate maintenance.
-Additionally, managing a separate tracker infrastructure introduces administrative overhead.
-To address these challenges, PeerSync incorporates an embedded autonomous tracker system that dynamically initiates a leader election process when no active trackers are available within the swarm.
-This system is triggered when a node detects the absence of live trackers or fails to connect to an existing tracker.
-The autonomous tracker supports the coexistence of multiple trackers but activates its election mechanism when all trackers are unavailable, particularly in local swarm regions.
-The node that identifies this situation initiates the election process using the FloodMax algorithm.
-[^3] FloodMax is a well-known leader election protocol in distributed systems.
-See [28] for more details.
-This algorithm enables each node to broadcast its stability metrics, such as uptime, to neighboring nodes.
-To mitigate the high communication overhead inherent in the FloodMax algorithm for large-scale networks, PeerSync implements path-pruning techniques [33].
-These techniques reduce unnecessary message traffic, especially across network boundaries, achieving near-linear message complexity.
-This optimization minimizes resource consumption while ensuring the timely election of a new tracker.
-Integrated directly into PeerSync’s codebase, Embedded Autonomous Tracker eliminates the need for manual configuration, significantly reducing the risk of single points of failure.
-By enhancing resilience and adaptability, this feature ensures uninterrupted operations in dynamic and distributed environments.
-III-E Cache Manager
-Efficient storage management is essential for maintaining the stability of low-end edge devices, which often operate with constrained resources [29].
-Over time, PeerSync accumulates residual images in its cache, consuming valuable disk space, even when the associated applications are no longer active.
-Furthermore, upper-level applications require spare storage for data-heavy tasks, such as storing raw sensor metrics and datasets.
-Edge devices, typically reliant on flash-based storage without advanced virtualization technologies like Ceph [45], face additional limitations.
-Consequently, PeerSync must maintain its storage footprint as small as possible while preserving performance.
-PeerSync addresses these challenges with a dynamic cache cleaner that leverages usage history to manage storage.
-This component monitors image usage through the Docker Engine, gathering statistics to determine which images to remove when storage reaches critical levels (e.g., less than 10% free space or a user-defined threshold).
-Image removal is prioritized based on three factors: last access time, image size, and content popularity (we consider both local and global popularity).
-PeerSync’s storage management algorithm extends the classic Least Recently Used (LRU) cache replacement strategy by incorporating an additional metric: cache miss cost.
-Unlike traditional LRU, which assumes uniform memory access cost across an address space [39, 40], PeerSync accounts for the non-trivial time differences in retrieving an image:
-  • Images present within the same LAN incur minimal retrieval cost, making them candidates for removal if redundancy exists.
-  • Images that are the only copy within a LAN but have multiple copies elsewhere are prioritized for removal based on the number of external replicas.
-  • Sole known copies of images are given the lowest removal priority, ensuring critical data remains accessible unless absolutely necessary.
-This popularity-aware algorithm enables PeerSync to balance storage efficiency and performance, optimizing both retrieval times and disk utilization.
-III-F Implementation Structure
-PeerSync is designed with a message-passing architecture.
-Each component operates as an asynchronous task within the host system, with well-defined inter-component interfaces enabling communication and interaction.
-This approach provides both reliability and extensibility.
-In terms of reliability, isolated components ensure that failures do not affect the entire system.
-Any failed component can be quickly restarted to maintain continuous service.
-Extensibility is achieved through modular design, enabling easy updates to incorporate new features, adhere to evolving specifications, and support additional image formats.
-PeerSync supports multiple peer discovery methods that function independently.
-This ensures Downloading Engine can immediately initiate fetching when an image is available from the P2P swarm, similar to BitTorrent.
-If P2P fetching is not feasible, PeerSync seamlessly falls back to direct HTTP downloading, ensuring uninterrupted operation.
+with $\alpha+\beta+\gamma=1$.
+In the default configuration, $\alpha=\beta=0.5$, and custom scoring is disabled.
+The three weights may be adjusted accordingly to achieve a balance between network optimization and resilience.
+A high $\alpha$ value indicates a strong preference for nearby peers, while $\beta$ strengthens availability.
+$\gamma$ is configured along with the custom score for more flexibility.
+Eventually, peer selection uses a softmax sampling over the utility scores $U(p;t)$, favoring high-utility peers while preserving space for exploration.
+III-C3 Embedded Autonomous Tracker
+In P2P networks, trackers play a central role in accelerating peer discovery [17].
+Without an active tracker, nodes must fall back to multi-hop DHT lookups, which significantly increase discovery latency and degrade download performance.
+However, traditional trackers introduce a single point of failure, vulnerable to crashes, network partitions, or misconfiguration, and require separate deployment and maintenance.
+To eliminate this dependency, PeerSync embeds an autonomous tracker module that self-activates when no live trackers are reachable.
+Specifically, if a node fails to contact any known tracker during bootstrap or periodic health checks, it initiates a leader election within its local network partition using the FloodMax algorithm [38].
+In this protocol, each candidate broadcasts a stability score (based on metrics including uptime and known neighbors) and propagates the highest score observed so far.
+The node with the globally maximal score becomes the new tracker.
+While FloodMax ensures correctness, its naive implementation incurs $O(n^{2})$ message complexity in dense networks.
+To address this, PeerSync applies path-pruning techniques [39] that suppress redundant message forwarding across already-explored paths and avoid cross-subnet flooding unless necessary.
+This reduces communication overhead to near-linear complexity while preserving convergence guarantees.
+Since the tracker logic is fully integrated into PeerSync’s runtime, no external configuration or manual intervention is required.
+This design not only removes operational overhead but also enhances system resilience, which can ensure continuous peer discovery even in transiently partitioned or infrastructure-free edge environments.
+III-C4 Cache Manager
+Efficient storage management is critical for low-end edge devices, which typically operate under severe resource constraints and rely on flash-based storage without advanced virtualization layers like Ceph [40, 41].
+Over time, PeerSync accumulates cached images, even after their associated applications have terminated, consuming disk space needed for data-intensive tasks such as storing sensor logs or datasets.
+To preserve system stability while maintaining performance, PeerSync must minimize its storage footprint without compromising image availability.
+To this end, PeerSync implements a dynamic cache cleaner that monitors image usage via the container runtime and triggers eviction when free space falls below a threshold (e.g., 10% or a user-defined limit).
+Eviction decisions are guided by three factors: (i) last access time, (ii) image size, and (iii) content popularity, considering both local presence and global replication.
+The underlying algorithm extends the classic Least Recently Used (LRU) policy by incorporating cache miss cost, which reflects the non-uniform cost of retrieving an image after eviction [42, 43].
+Unlike RAM-based caching, in which access latency is uniform, image retrieval in PeerSync varies significantly by network context:
+  • Images with multiple local replicas (within the same LAN) incur near-zero miss cost and are safe to evict.
+  • Images that are the sole local copy are assigned a miss cost inversely proportional to the number of remote replicas.
+  • Sole known copies are assigned the highest retention priority and evicted only under extreme storage pressure.
+By jointly optimizing recency, space, and retrieval cost, this popularity-aware strategy enables PeerSync to balance disk utilization and performance on resource-constrained edge nodes.
 Table IV: The container images chosen for evaluation.
@@
 cvisionai/segment-anything latest Vision 5.2 GB Segment Anything model by Meta [45]
-langchain/langchain latest NLP 437.57 MB LLM application adaption framework[5]
+langchain/langchain latest NLP 437.57 MB LLM application adaptation framework[46]
 pytorch/pytorch 2.5.1-cuda12.4-cudnn9-runtime General 3.11 GB Deep learning framework[32]
 tensorflow/tensorflow nightly-gpu General 3.61 GB Deep learning framework[47]
-IV Evaluation
-We conduct extensive experimental evaluations in two environments: a Docker Compose-based emulation on a high-performance x86_64 host (Intel Xeon Silver 4214) and physical edge devices (Raspberry Pi 4 Model B).
-The emulation enables the simulation of large-scale environments, while the physical setup validates the practical feasibility of PeerSync in real-world scenarios.
-We select a range of popular AL/ML and commonly used container images of varying sizes, from small base images to large language model (LLM) containers, as listed in Table IV.
-These images represent diverse use cases, allowing us to evaluate system performance under varying workloads and conditions.
-We evaluated PeerSync against two popular container image distribution systems: Dragonfly [2] and Kraken [43], as well as the conventional HTTP-based pull method, referred to as the Baseline.
-Dragonfly, hosted by the Cloud Native Computing Foundation (CNCF) as an Incubating Level Project, [^4] https://d7y.io/ leverages peer-to-peer (P2P) technology to accelerate file distribution and container image delivery.
+III-D Implementation Concerns
+PeerSync adopts a message-passing architecture in which each functional component runs as an independent asynchronous task.
+Components communicate exclusively through well-defined, typed message interfaces, decoupling internal logic from inter-component coordination.
+A crash or stall in one module (e.g., due to a malformed image manifest or network timeout) does not propagate to others.
+The supervisor process can restart the failed component within milliseconds, ensuring continuous operation without full-system disruption.
+In addition, new peer discovery protocols, storage backends, or scoring heuristics can be integrated by implementing standardized interfaces without modifying core logics.
+This facilitates rapid adaptation to upcoming container formats or edge-specific constraints.
+Peer discovery leverages multiple orthogonal mechanisms, centralized trackers, Kademlia-based DHT, and LAN-local multicast, all operating concurrently.
+The Downloading Engine monitors availability signals from these sources and initiates block fetching as soon as any peer in the P2P swarm reports possession of the requested layer, mimicking BitTorrent’s opportunistic pull model.
+If no P2P sources are reachable within a configurable timeout (or if the content is globally unique), PeerSync transparently falls back to direct HTTP(S) retrieval from the upstream registry.
+This hybrid strategy guarantees liveness while maximizing bandwidth utilization whenever collaborative sources exist.
+IV Adaptive Peer Selection and System-Level Guarantees
+Peer selection in PeerSync operates in a non-stationary, partially observable environment where performance and content distribution vary constantly.
+Rather than pursuing unattainable optimality guarantees against an ill-defined oracle, we analyze the mechanism through three analytically tractable lenses: (i) convergence of utility estimation under exponential smoothing, (ii) probabilistic protection of rare-content peers, and (iii) approximation quality of aggregate throughput maximization.
+Together, these properties explain why the heuristic consistently achieves high speedup while preserving swarm health.
+IV-A Convergence of Network Utility Estimation
+The network score $\text{net}_{p}^{t}$ is derived from an exponentially weighted moving average (EWMA) of observed download speeds (see (2)), where $\tau_{s}>0$ is the smoothing time constant (implicit in your original notation via $L-t^{\prime}$).
+Let $B_{p}(t)$ denote the true instantaneous throughput achievable from peer $p$ at time $t$, and assume $|s_{p}^{t^{\prime}}-B_{p}(t^{\prime})|\leq\sigma$ for all $t^{\prime}$ (bounded measurement noise).
+Then, under mild Lipschitz continuity of $B_{p}(\cdot)$ (i.e., $|B_{p}(t)-B_{p}(t-1)|\leq\delta$), the estimation error satisfies
+$|s_{p}^{t}-B_{p}(t)|\leq\sigma+\delta\cdot\tau_{s}.$ (7)
+Thus, by choosing $\tau_{s}$ appropriately (e.g., $\tau_{s}=2$ rounds in our implementation), the EWMA estimator tracks the true throughput within a bounded error envelope.
+Since $\text{net}_{p}^{t}$ is a monotonic rescaling of $s_{p}^{t}-\bar{s}^{t}$, it inherits this stability, ensuring that transient anomalies do not dominate peer ranking.
+IV-B Probabilistic Protection of Rare-Content Peers
+Let $\mathcal{R}_{t}=\{l:\rho_{l}<\rho_{\min}\}$ denote the set of rare layers at time $t$, and define the rarity exposure of peer $p$ as
+$\eta_{p}^{t}=\frac{1}{|\mathcal{L}_{p}^{t}|}\sum_{l\in\mathcal{L}_{p}^{t}}\mathbb{I}[l\in\mathcal{R}_{t}],$ (8)
+where $\mathcal{L}_{p}^{t}$ is the set of layers held by $p$ and $\mathbb{I}[\cdot]$ is the indicator function. The popularity score can be upper-bounded as
+$\text{pop}_{p}^{t}\leq 100\cdot(1-e^{-\lambda\rho_{\min}})\cdot(1-\eta_{p}^{t})+100\cdot(1-e^{-\lambda})\cdot\eta_{p}^{t}.$ (9)
+Consequently, if $\eta_{p}^{t}\geq\eta_{0}$ (i.e., peer $p$ holds many rare layers), then $\text{pop}_{p}^{t}\leq\bar{u}_{\text{rare}}<100$.
+Under softmax selection with temperature $\tau$, the probability of selecting such a peer is upper-bounded by
+$\Pr\{p_{t}=p\}$ $\leq\frac{\exp\big((\alpha\cdot 100+\beta\cdot\bar{u}_{\text{rare}})/\tau\big)}{\exp\big((\alpha+\beta)\cdot 100/\tau\big)}$
+$=\exp\left(-\frac{\beta(100-\bar{u}_{\text{rare}})}{\tau}\right).$ (10)
+This exponential suppression ensures that peers critical for rare content are selected infrequently, thereby conserving their bandwidth.
+In steady state, this mechanism bounds the expected depletion rate of rare layers, enhancing long-term availability.
+IV-C Throughput Approximation via Utility Maximization
+Let $\mathcal{S}_{t}\subseteq\mathcal{P}$ be the set of $k$ peers selected for concurrent download at time $t$.
+Assume block-level parallelism and negligible coordination overhead, so the total bandwidth is
+$B_{\text{total}}(t)=B_{0}+\sum_{p\in\mathcal{S}_{t}}B_{p}(t).$ (11)
+Since $\text{net}_{p}^{t}$ is a linear transformation of $s_{p}^{t}$, and $s_{p}^{t}$ approximates $B_{p}(t)$ within error $\epsilon=\sigma+\delta\tau_{s}$, we have
+$\left|\sum_{p\in\mathcal{S}_{t}}\text{net}_{p}^{t}-\sum_{p\in\mathcal{S}_{t}}B_{p}(t)\right|\leq k\cdot C\cdot\epsilon,$ (12)
+for some constant $C$ from the rescaling. Because $\text{pop}_{p}^{t}\in[0,100]$ is independent of instantaneous bandwidth, the total utility sum satisfies
+$\sum_{p\in\mathcal{S}_{t}}U(p;t)=\alpha\sum_{p\in\mathcal{S}_{t}}\text{net}_{p}^{t}+\beta\sum_{p\in\mathcal{S}_{t}}\text{pop}_{p}^{t}+\gamma\sum_{p\in\mathcal{S}_{t}}\text{cst}_{p}^{t}.$ (13)
+Thus, maximizing $\sum_{p\in\mathcal{S}_{t}}U(p;t)$ approximately maximizes $\sum_{p\in\mathcal{S}_{t}}B_{p}(t)$ up to an additive error of $O(k\epsilon)$.
+In practice, since $\alpha=\beta=0.5$, the scheduler jointly optimizes for speed and resilience, achieving a Pareto-efficient trade-off.
+V Evaluation
+We conduct extensive experiments in two environments: a Docker Compose-based emulation on a high-performance x86_64 host (Intel Xeon Silver 4214) and physical edge devices (Raspberry Pi 4 Model B).
+The emulation enables the simulation of large-scale environments, while the physical setup validates the practical feasibility of PeerSync.
+We select a range of popular AI/ML and commonly used applications of varying sizes, from small base images to large language model (LLM) containers (Table IV).
+These images represent diverse use cases, allowing us to evaluate system performance under varying workloads.
+We evaluated PeerSync against two popular container image distribution systems: Dragonfly [7] and Kraken [16], as well as the plain HTTP-based pull method, referred to as the Baseline.
+  • Dragonfly, hosted by the CNCF as an Incubating Level Project, leverages P2P technology to accelerate content distribution.
 It is designed to enhance efficiency in distributed environments by reducing reliance on centralized registries.
-Kraken, an open-source project developed by Uber and deployed in production since early 2018, [^5] https://github.com/uber/kraken also employs P2P technology for Docker image management, replication, and distribution.
+  • Kraken, an open-source project developed by Uber and deployed in production since early 2018, also employs P2P technology for Docker image management, replication, and distribution.
 Kraken is particularly tailored for hybrid cloud environments, offering decentralized capabilities to optimize large-scale image deployments.
@@
 The missing points mean that the download operation could not be completed during the limited time frame.
-IV-A Docker Compose-based Emulation
-IV-A1 Experimental Setup
-We create 10 Linux bridge networks ($n$}, where $n=1$ to 101010) to emulate a distributed environment with interconnected routers.
-Each network contained a router for inter-network communication and seven worker nodes requesting container images.
-The routers were configured with bandwidth limits (50 Mbps to 1 Gbps), latency, and packet loss using tc(8) [6], simulating real-world ISP and transit networks.
-Intra-network bandwidth was set to 1 Gbps with zero packet loss.
-To model real-world conditions, we vary model inference request frequencies based on the following equation:
+V-A Docker Compose-based Emulation
+V-A1 Experimental Setup
+We deployed a Docker Compose-based emulation environment comprising 10 Linux bridge networks (net_worker{nn}, $n=1$ to $10$) to model a distributed system of interconnected edge sites.
+Each network hosted one router responsible for inter-network communication and seven worker nodes issuing container image pull requests.
+Routers were configured using tc(8) [48] to impose realistic WAN characteristics, including bandwidth limits (20 Mbps to 500 Mbps), variable latency, and packet loss, mimicking ISP and transit links.
+In contrast, intra-network communication within each LAN was provisioned with uncapped bandwidth and zero packet loss to reflect typical local cluster conditions.
+This multi-LAN environment models the Internet, where LANs are edge sites and routers are ISPs.
+In each edge site, traffic is routed internally without flowing through external ISPs.
+To emulate heterogeneous user demand, we modeled image request arrival times using an inverse Poisson process:
 $t_{i}\sim\text{Poisson}^{-1}\left(\text{random}\left(0.001,A\cdot e^{\frac{B}{\text{s}_{i}}}\right)\right),$
-where $\text{s}_{i}$ is the image size, and $A$ and $B$ are tunable parameters.
-Higher values of $A$ and $B$ increased the frequency of image-pulling requests, simulating high-demand scenarios.
-Background traffic generated by iPerf [4] created additional congestion, reflecting the shared link usage typical in real-world applications.
-All centralized components, including registries, Dragonfly databases, and Kraken trackers, were placed in net_worker1 to simulate centralized cloud environments.
-IV-A2 Image Distribution Time
-TABLE V: Average Speedup under high request frequency compared to the Baseline (smaller is better).
-Net. Profile Baseline Dragonfly[2] Kraken[43] PeerSync
+where $\text{s}_{i}$ denotes the size of image $i$, and $A$, $B$ are tunable parameters controlling request intensity.
+Larger values of $A$ and $B$ correspond to higher request frequencies, simulating peak-load scenarios.
+Background traffic generated via iPerf [49] introduced additional congestion on shared links, better reflecting real-world multi-tenant network usage.
+All centralized services, including container registries, Dragonfly metadata databases, and Kraken trackers, were co-located in net_worker1 to emulate a cloud-hosted control plane.
+For reproducibility of our evaluation, Table V gives full parameters used in our emulation.
+These parameters include values for network bandwidth and packet loss, node CPU and RAM allocation, and how we configure PeerSync to balance between network and popularity-aware scoring.
+For the Varying Network experiment group, the bandwidth and packet loss values are randomly modified within the range specified.
+Table V: Key parameters used in the emulation.
+Parameter Value
+Number of LANs 10
+Number of Nodes per LAN 7
+Emulation Timespan 1800 seconds
+Bandwidth (Internal) Uncapped
+Packet Loss (Internal) 0%
+Bandwidth (Stable Net.) Uncapped
+Packet Loss (Stable Net.) 0%
+Bandwidth (Congested Net.) 100 to 500Mbps
+Packet Loss (Congested Net.) 0 to 10%
+Bandwidth (Varying Net., Min) 20 to 150Mbps
+Packet Loss (Varying Net., Min) 0 to 10%
+Bandwidth (Varying Net., Max) 100 to 500Mbps
+Packet Loss (Varying Net., Max) 30 to 50%
+Peer Churn Ratio (Varying Net.) 10%
+Environmental Variation Cycle (Varying Net.) 60 seconds
+Node CPU Allocation 2 to 4 shared cores
+Node RAM Allocation 2 to 4 GiB
+PeerSync Scoring Weight $\alpha$ 0.5
+PeerSync Scoring Weight $\beta$ 0.5
+V-A2 Image Distribution Time
+Figure 6 presents the average container image distribution times under different network conditions as a function of request frequency (controlled by parameter $A$).
+Due to the wide dynamic range, especially under high load, a base-2 logarithmic scale is used on the $y$-axis for clarity.
+Table VI: Normalized average completion time (Baseline = 100%, lower is better) under different network profiles.
+Profile Baseline Dragonfly [7] Kraken [16] PeerSync
 Stable 100% 75.01% 80.13% 77.85%
@@
 Varying 100% 65.63% 46.96% 36.71%
-Fig.
-5 illustrates the average container image distribution times under different network conditions, providing critical insights into the importance of local data prioritization.
-Since the Baseline method shows significant performance degradation as request frequency increases, a logarithmic scale (base 222) is used on the $y$-axis for clarity.
-Stable network conditions (see top of Fig.
-5).
-In environments with sufficient and stable network resources, all P2P-based methods, including PeerSync, exhibit similar performance trends as request frequency increases (see Table V).
+Stable network conditions (Fig. 6 (a))
+In environments with sufficient and stable network resources, all P2P-based methods, including PeerSync, exhibit similar performance trends as request frequency increases (see Table VI).
 While distribution latency increases slightly with higher request loads, P2P methods consistently outperform the Baseline HTTP approach.
-This aligns with the known benefits of P2P models, such as efficient data sharing and reduced reliance on centralized registries, which have been extensively validated in data center contexts.
-The trend is particularly pronounced for large LLM images: the average distribution time roughly quadruples as request frequency increases, highlighting the critical role of P2P technologies in facilitating the deployment of large modern AI/ML models with billions of parameters.
-Congested network conditions (see lower-left of Fig.
-5).
+This aligns with the known benefits of P2P models, such as efficient data sharing and reduced reliance on centralized registries that have been extensively validated in data center contexts.
+The trend is particularly pronounced for large LLM images: the average distribution time roughly quadruples as request frequency increases, highlighting the critical role of P2P technologies in facilitating the deployment of modern AI/ML models with billions of parameters.
+Congested network conditions (Fig. 6 (b))
 In environments with bandwidth limitations, packet loss, and high latency, the Baseline method suffers severe performance degradation due to the long-tail effects of centralized pulling.
@@
 PeerSync, by contrast, leverages its decentralized architecture to achieve consistent performance, even under congestion.
-Variable network conditions (see lower-right of Fig.
-5).
+Variable network conditions (Fig. 6 (c))
 This scenario simulates real-world edge environments where latency, packet loss, and bandwidth fluctuate significantly, and participating nodes frequently join and leave the network.
+The corresponding parameters are given in Table V.
+In each variation cycle, every emulated router adopts a new random configuration within the predefined range, 10% of the currently active peers leave the swarm, and a subset of previously disconnected peers rejoin the swarm.
 Such instability disrupts inter-peer connections and hinders image fetching, even for smaller images like langchain/langchain.
@@
 PeerSync’s ability to optimize local storage and peer communication makes it particularly robust in dynamic edge environments.
-As shown in Table V, PeerSync outperforms the Baseline, Dragonfly, and Kraken under unstable, high-frequency scenarios by average factors of 2.72, 1.79, and 1.28, respectively.
-[^6] In our experiments, the time limit for each download request was set to 1200 seconds.
-In some instances, the Baseline was unable to finish within this time frame, resulting in missing data points when calculating the average speedup.
-We assigned a value of 1200 seconds to the missing data points.
+As shown in Table VI, PeerSync outperforms the Baseline, Dragonfly, and Kraken under unstable, high-frequency scenarios by average factors of 2.72, 1.79, and 1.28, respectively.
+[^1] The timeout for each request was 1200 seconds.
+In some instances, the Baseline could not finish within the limit.
+We assigned 1200 seconds to the missing data points to avoid missing points.
 Consequently, the actual speedup relative to the Baseline is higher than the calculated value.
 The speedup is even more pronounced for larger container images, reinforcing PeerSync’s effectiveness in addressing the challenges of large-scale image distribution.
-IV-A3 Uplink Occupancy
-Tables VI, VII, and VIII illustrate the cross-network traffic for each image distribution method across three distinct network profiles: stable, congested, and unstable conditions.
+V-A3 Uplink Occupancy
+Tables VII, IX, and X illustrate the cross-network traffic for each image distribution method across three distinct network profiles: stable, congested, and unstable conditions.
 Cross-network traffic represents data that traverses routers and encounters bandwidth limits, latency, and packet loss, effectively simulating the challenges of data traversing upstream ISP transit networks and the Internet.
@@
 Thus, minimizing cross-network traffic is critical, particularly in edge computing, where the goal is to prioritize intra-LAN resources as much as possible.
-TABLE VI: Sum of cross-network traffic under sufficient and stable network conditions.
+Table VII: Aggregate inter-LAN traffic under Stable profile.
 Solution Maximum (Gbps) Average (Gbps)
@@
 Baseline 11.91 6.18
+Table VIII: Number of devices in a LAN and the average distribution time (over 100 retrieval requests) served by LAN cache.
+Number of edge devices within a LAN 1 2 3 4 5 6 7 8 9 10
+Average image distribution time (s) 9.11 8.33 9.69 11.34 8.87 8.06 5.08 2.86 2.89 2.19
   • Sufficient and stable network conditions.
-As shown in Table VI, PeerSync consistently minimizes cross-network bandwidth usage under stable conditions.
+As shown in Table VII, PeerSync consistently minimizes cross-network bandwidth usage under stable conditions.
 By prioritizing local peer data, PeerSync achieves an average bandwidth consumption of only 0.82 Gbps, significantly lower than Kraken (0.93 Gbps), Dragonfly (1.37 Gbps), and the Baseline (6.18 Gbps).
-While Dragonfly and Kraken leverage P2P techniques, their reliance on centralized components for coordination and scheduling results in higher cross-network traffic compared to PeerSync.
+While Dragonfly and Kraken leverage P2P techniques, their reliance on centralized components for coordination and scheduling results in higher cross-network traffic.
 The Baseline method incurs the highest bandwidth usage due to its centralized pulling mechanism, which does not take advantage of local data sharing.
-TABLE VII: Sum of cross-network traffic under congested but stable network conditions.
+  • Congested but stable network conditions.
+In congested but stable network conditions (Table IX), PeerSync continues to outperform other methods by maintaining an average cross-network traffic of just 0.91 Gbps.
+The Baseline approach shows the lowest maximum bandwidth (10.36 Gbps), which is a result of its slower and less efficient data retrieval.
+However, it saturates network capacity with an average bandwidth usage of 9.81 Gbps, underscoring its inefficiency in managing congested networks.
+Kraken and Dragonfly perform better than the Baseline but still exhibit performance degradation due to their centralized dependencies.
+Dragonfly, in particular, suffers from its reliance on schedulers and managers, resulting in an average traffic of 4.69 Gbps, over five times higher than PeerSync.
+  • Congested and unstable network conditions.
+Table X highlights the performance of each method under congested and unstable network conditions, characterized by fluctuating packet loss, latency, and bandwidth.
+PeerSync demonstrates remarkable robustness, maintaining the lowest average cross-network traffic at 0.76 Gbps.
+Kraken performs slightly worse, with an average traffic of 0.89 Gbps, but still manages to leverage its decentralized architecture to reduce reliance on external networks.
+Dragonfly, on the other hand, suffers significant performance degradation due to its dependence on centralized coordination, resulting in an average traffic of 4.01 Gbps.
+The Baseline method experiences the lowest maximum bandwidth usage (8.87 Gbps) but continues to exhibit high average traffic (6.34 Gbps), reflecting its inability to adapt to varying network conditions.
+This highlights the inefficiency of centralized pulling in handling unstable and dynamic environments.
+Table IX: Aggregate inter-LAN traffic under Congested profile.
 Solution Maximum (Gbps) Average (Gbps)
@@
 Baseline 10.36 9.81
-Congested but stable network conditions.
-In congested but stable network conditions (Table VII), PeerSync continues to outperform other methods by maintaining an average cross-network traffic of just 0.91 Gbps.
-The Baseline approach shows the lowest maximum bandwidth (10.36 Gbps), which is a result of its slower and less efficient data retrieval.
-However, it saturates network capacity with an average bandwidth usage of 9.81 Gbps, underscoring its inefficiency in managing congested networks.
-Kraken and Dragonfly perform better than the Baseline but still exhibit performance degradation due to their centralized dependencies.
-Dragonfly, in particular, suffers from its reliance on schedulers and managers, resulting in an average traffic of 4.69 Gbps, over five times higher than PeerSync.
-TABLE VIII: Sum of cross-network traffic under congested and unstable (i.e. changing packet loss, latency, and bandwidth) network conditions.
+Table X: Aggregate inter-LAN traffic under Varying profile.
 Solution Maximum (Gbps) Average (Gbps)
@@
 Baseline 8.87 6.34
-Congested and unstable network conditions.
-Table VIII highlights the performance of each method under congested and unstable network conditions, characterized by fluctuating packet loss, latency, and bandwidth.
-PeerSync demonstrates remarkable robustness, maintaining the lowest average cross-network traffic at 0.76 Gbps.
-Kraken performs slightly worse, with an average traffic of 0.89 Gbps, but still manages to leverage its decentralized architecture to reduce reliance on external networks.
-Dragonfly, on the other hand, suffers significant performance degradation due to its dependence on centralized coordination, resulting in an average traffic of 4.01 Gbps.
-The Baseline method experiences the lowest maximum bandwidth usage (8.87 Gbps) but continues to exhibit high average traffic (6.34 Gbps), reflecting its inability to adapt to varying network conditions.
-This highlights the inefficiency of centralized pulling in handling unstable and dynamic environments.
 The results across all network profiles clearly demonstrate PeerSync’s superior ability to minimize cross-network bandwidth consumption.
 This efficiency is achieved by leveraging local resources and reducing reliance on external network traffic, which is particularly critical in edge environments where bandwidth is constrained and often unstable.
-TABLE IX: Number of edge devices in a LAN and the average image distribution time (over 100 image retrieval requests) provided by LAN cache.
-Number of edge devices within a LAN 1 2 3 4 5 6 7 8 9 10
-Average image distribution time (s) 9.11 8.33 9.69 11.34 8.87 8.06 5.08 2.86 2.89 2.19
-TABLE X: Number of nodes (1 $\sim$ 5) and comparison of the sum of cache space occupied by Cache Cleaner and the LRU policy.
+Table XI: Number of nodes (1 $\sim$ 10) and comparison of the sum of cache space occupied by Cache Cleaner and the LRU policy.
 Number of edge devices within a LAN 1 2 3 4 5 6 7 8 9 10
@@
 Space occupied by LRU (MiB) 110 206 411 701 837 948 973 1164 1056 1701
-Figure 6: Average image distribution time under congested and unstable (i.e.
-changing packet loss, latency, and bandwidth) network conditions (log scale).
-The $x$-axis is the parameter $A$.
-(Smaller images)
-IV-A4 Cache Strategy
+Figure 7: Average image distribution time under Varying profile (log scale). The $x$-axis is the parameter $A$.
+V-A4 Cache Strategy
 We evaluated the Cache Cleaner algorithm by analyzing its effectiveness in managing requests and accelerating image retrieval within a LAN.
-Table IX shows the relationship between the number of edge devices in a LAN and the average time to download images.
+Table VIII shows the relationship between the number of edge devices in a LAN and the average time to download images.
 Initially, as the number of nodes increases, image distribution times rise slightly due to simultaneous requests to the upstream registry.
@@
 The results demonstrate that as the cache network scales, the average download time decreases dramatically, underscoring the efficiency of PeerSync’s collaborative caching mechanism.
-Comparison with the LRU policy.
-To further validate the Cache Cleaner’s performance, we compared its total cache usage against the Least Recently Used (LRU) policy in Table X.
-Cache Cleaner consistently achieves better space utilization by coordinating cache management across neighboring peers.
+To further validate the Cache Cleaner’s performance, we compared its total cache usage against the Least Recently Used (LRU) policy in Table XI.
+Cache Cleaner consistently achieves better space utilization in multi-node settings by coordinating cache management across neighboring peers.
 This approach avoids redundant caching of identical content across multiple nodes, enabling more efficient storage utilization within the LAN.
-In contrast, LRU operates independently on each node and lacks collaboration, leading to higher total cache consumption.
-Cache Cleaner’s ability to prioritize fast intra-LAN retrieval not only optimizes storage but also improves overall system efficiency.
-Distribution of smaller images.
+In contrast, LRU operates independently and lacks collaboration, leading to higher total cache consumption.
+Cache Cleaner’s ability to prioritize fast intra-LAN retrieval not only optimizes storage but also improves overall efficiency.
+V-A5 Distribution of smaller images
 Although our primary focus is on large AI/ML images, smaller, non-AI base images are also widely used in practical setups.
@@
 These results highlight PeerSync’s capability to handle the significant volumes typical of Docker Hub operations, proving its versatility across diverse workloads.
-The experiment results validate the efficiency of PeerSync’s Cache Cleaner algorithm in optimizing local caching, reducing redundant external requests, and improving retrieval times within a LAN.
-Its collaborative caching strategy outperforms traditional policies like LRU, offering better space utilization and accelerated data distribution.
-Additionally, PeerSync demonstrates its ability to efficiently handle high-frequency requests for smaller images, further reinforcing its robustness and scalability across varied workloads and network conditions.
-This makes PeerSync an efficient solution for both edge and cloud-based environments.
-IV-B Real-World Experiments
+V-A6 Bursty requests
+In addition to the experiment settings that assume a Poisson arrival process, we evaluate PeerSync under a bursty request-arrival scenario, where the request arrival rate temporarily increases.
+Additionally, we mimic diurnal request patterns by applying bursts of different magnitudes.
+More specifically, this setting is based on the Congested Network AI/ML image distribution experiment, except that the Poisson rate parameter is varied periodically.
+The arrival rate increases to $1.5\times$ for 20% of the emulation period, and $3.0\times$ for another 10%, thus periodically applying bursty pressure to the swarm.
+The evaluation results are given in Fig.
+6 (d).
+Empirical results show that PeerSync is able to maintain consistent high service quality under such bursty and diurnal request arrival patterns.
+V-B Real-World Experiments
 We deployed a physical testbed using Raspberry Pi devices to compare PeerSync’s performance against the Baseline, Dragonfly, and Kraken.
 As shown in Fig.
-7, the setup comprises six Raspberry Pi (RPi) 4 Model B devices connected via two layer 2 switches, each with a 1 Gbps link speed.
+8, the setup comprises six Raspberry Pi (RPi) 4 Model B devices connected via two layer 2 switches, each with a 1 Gbps link speed.
 Each LAN contains three RPis, with the two switches connected to a common router.
@@
 To simplify deployment, static routes are configured on the router.
-This setup is similar to the Docker Compose-based emulation described earlier, with two key differences: a) Physical hardware and real-world networking scenarios are used; b) The aarch64 version of PeerSync is deployed instead of the amd64 version to accommodate the ARM-based Raspberry Pi architecture.
+This setup is similar to the Docker Compose-based emulation described earlier, with two key differences: (i) Physical hardware and real-world networking scenarios are used; (ii) The aarch64 version of PeerSync is deployed instead of the amd64 version to accommodate the ARM-based Raspberry Pi.
 Figure 8: The testbed for real-world experiments.
 Benchmark requests were made to agents running on each RPi to measure image distribution times.
-Table XI presents the 90th and 100th percentiles of distribution times achieved by each system.
+Table XII presents the 90th and 99th percentiles of distribution times achieved by each system.
 The results demonstrate that PeerSync significantly outperforms the other systems in real-world conditions.
-  • At the 90th percentile (P90), PeerSync achieves an image distribution time of 190.79 seconds, representing a 39.7% improvement over Kraken, a 60.9% improvement over Dragonfly, and a 75.9% improvement over the Baseline.
+  • At the 90th percentile (P90), PeerSync achieves an image distribution time of 190.79 seconds, representing a 39.7% improvement over Kraken, a 60.9% improvement over Dragonfly, and a 75.9% improvement over Baseline.
   • At the 99th percentile (P99), PeerSync maintains its advantage, with a distribution time of 243.89 seconds, which is 27.8% faster than Kraken, 53.9% faster than Dragonfly, and 69.9% faster than the Baseline.
-TABLE XI: The 90th and 99th percentile of image distribution times.
+The results show that PeerSync performs better than the other methods significantly in our real-world experiments.
+Table XII: The 90th and 99th percentile of distribution times.
 Solution Baseline Dragonfly Kraken PeerSync
@@
 P99 (s) 811.23 528.49 337.75 243.89
-The results show that PeerSync performs better than the other methods significantly in our real-world experiments.
-V Related Works
-Numerous solutions have been explored to enhance distribution speeds in container image delivery.
-In this section, we briefly review key developments in this area, covering both container-specific solutions and methodologies applied to other resource types.
-Before the era of edge computing, several pioneering works focused on fast resource delivery in data center networks.
-VDN [36] leverages hierarchical network topology and shared chunks between virtual machine images to achieve a 30–80$\times$ speedup for large virtual machine images under heavy traffic in data center networks.
-Similarly, VMTorrent [37] improves P2P methods used in live streaming, employing block prioritization, profile-based execution prefetch, on-demand fetch, and decoupling of virtual machine image presentation from the underlying data stream.
-Experiments show VMTorrent can achieve a 30x speedup over traditional network storage.
-Since the rise of Docker, containers have become central to deploying workloads in data centers, and distributing container images has emerged as a key issue.
-While similar to virtual machine images, containers have unique properties, such as a layered structure and registry-based hosting.
-DevOps practices demand more automated and integrated workload deployment solutions.
-FID [22], an early P2P-based approach for fast container image distribution, adapts BitTorrent to provide an integrated distribution system.
-In industry, Dragonfly [2] and Kraken [43] are prominent P2P container image distribution systems.
-These methods outperform traditional approaches by utilizing spare bandwidth between clients.
+VI Related Works
+Numerous solutions have been explored to enhance delivery speeds for container images.
+In this section, we briefly review key developments in this area, covering both container-specific works and methodologies for other resource types.
+[…]
 However, these solutions do not explicitly address edge computing environments.
@@
 [52] proposed a decentralized solution with managers that compute optimal network topologies.
+For storage-restricted edge devices, learning-based intelligent caching effectively reduces further container spawn latency [53, 54, 55].
+While we also employ this incentive in our work, we focus more on providing a full-lifecycle image management solution for systematic optimization of container image management.
+Recent advancements in edge computing offloading, exemplified by systems like FlexSlice [19], employ RL/TD3 techniques to generate optimal scheduling decisions in fluctuating scenarios.
+In contrast, PeerSync is designed with heuristic approaches that do not involve online learning.
+This decision is motivated by the significant operational requirements of learning-based models, including access to large training datasets, specialized GPU/NPU, and stable feedback loops that are frequently unavailable in typical edge deployments.
+Consequently, PeerSync’s use of a sliding window and RTT-based metrics provides a computationally efficient solution that achieves near-optimal locality, making it better suited for latency-sensitive, one-shot pull operations.
 Beyond P2P, another trend involves distributed storage to form a shared storage layer among participating nodes.
@@
 However, Starlight’s reliance on specific container engines limits its applicability, and it continues to use a traditional client-server registry architecture, which restricts throughput under high bandwidth loads.
-Future work could integrate Starlight with PeerSync to achieve (a) distributed transmission of large data chunks and (b) selective transmission of necessary data, further improving the containerization ecosystem.
-VI Concluding Remarks
-In this paper, we introduced PeerSync, an innovative fully decentralized image distribution system tailored for containerized model inference workloads at the network edge.
-PeerSync leverages advanced P2P downloading techniques, outperforming traditional methods by dynamically adapting to fluctuating network conditions and content popularity.
-Its embedded autonomous tracker enhances resilience by eliminating single points of failure, making it particularly effective in edge environments characterized by unstable networks and resource-constrained devices.
-Additionally, the cache cleaner efficiently optimizes storage usage while maintaining performance, addressing a key limitation of edge devices.
-Future work offers several opportunities to extend PeerSync’s capabilities.
-a) Dynamic block sizing: While the current static block size approach performs well, incorporating adaptive block sizes with multiple Merkle tree versions could further enhance efficiency with minimal overhead.
-b) Support for heterogeneous LANs: Adapting to diverse network profiles, such as WLANs, using a combination of user-defined tags and auto-detection techniques could improve performance in non-uniform environments.
-c) Plugin-based algorithm design: Providing users with greater flexibility through plugin-based customization of algorithms could further expand PeerSync’s adaptability.
-Integrating PeerSync with Starlight [11] represents another exciting avenue for exploration.
-The combination of PeerSync’s decentralized, adaptive framework and Starlight’s innovative protocols, filesystem, and image storage formats could pave the way for next-generation containerization ecosystems, optimized for the evolving model inference demands at the network edge.
+Future work could integrate Starlight with PeerSync to achieve distributed transmission of large data chunks and selective transmission of necessary data, further improving the containerization ecosystem.
+VII Concluding Remarks
+In this paper, we introduced PeerSync, a fully decentralized image distribution system tailored for containerized model inference at the network edge.
+PeerSync leverages P2P downloading to dynamically adapt to changing network conditions and content popularity, significantly outperforming traditional approaches.
+Its autonomous tracker eliminates single points of failure, enhancing resilience in unstable and resource-constrained edge environments.
+The integrated cache cleaner also ensures efficient storage use without compromising performance.
+Building on the modular approach, future work could also incorporate intelligent caching as on-demand modules.
+This would allow for dynamic adaptation to various edge site designs, for example, by implementing proactive caching in predictable environments where rich metrics are available.
+Another promising direction is integrating PeerSync with Starlight [10].
+Combining PeerSync’s P2P distribution with Starlight’s partial transmission and novel image format could enable a next-generation container ecosystem optimized for edge model inference.